A Complete Bibliography of Publications in Reviews of Modern Physics (1950–1959)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.07

Title word cross-reference

(d^{2}ϕ/dx^{2}) + [a + b(e^{−x}/x)]ϕ = 0 [HL51]. 10 [FB56]. 11 [AH50a]. 1f_{7/2} [Nus56]. 20 [AH50a]. 30 [FB56]. 3d [WD58]. ^{++} [Hen53]. ^{+++} [Hen53]. ^{1}Σ [Whi55]. ^{∘} [RSB53]. _{2} [Gor53d, Hag51, WW58]. _{3} [HP53]. A [MMN51, Nor51]. β [sW50]. · [Gor53d]. ft [Fei51]. θ [AN59]. Z [AH50a]. Z ≤ 60 [Che56].

-MeV [FB56]. -Ray [sW50].

1-3 [Ano59b]. 1/2 [FV58]. 16-22 [Ano59d]. 1950 [Mac50]. 1955 [Bet55]. 1957 [Hof58]. 1958 [Ano59f].

21-cm [Ker58]. 26-31 [Ano59e].

32-35 [Ano59a]. 36-40 [Ano59g].

4-15 [Ano59i]. 41-46 [Ano59c]. 49-51 [Ano59h].

52-58 [Ano59j]. 59-61 [Ano59k].

=0 [Gor53b]. =11 [EK54, EB57]. =20 [EK54, EB57].

above [All58, All59b]. Absorption [DE52, HP53, KP52, Kip53, MBL51, Ric55, Ver59]. Abundance [AH50b, AH50c]. Abundances [SU56]. Accelerated [ABH^{+}56, ABH^{+}58]. Acetate [Ble53a]. Acids [Ano59e, Kor59, Lev59b, Pol59, Ric59c, Wil59b]. Acoustic [Bor53]. Action [Alb59, DeW57a, Inf57, Neu59]. Actions [Zir59]. Activity [Ros59b]. Adapted [Mel56]. Address [Whi58]. Adiabatic [dKH53]. Adjustment [CDLR55, DC53]. Aerodynamic [Pet58]. After [WvdH53]. After-Effect [WvdH53]. Agitation [Cri58, Lav58]. Air [Bet50]. Al [WW58]. Alkali [Sei54]. Alloy [SB58]. Alloys [BT58a, Gei53a, Gol53, Gor53a, NH53]. Alpha [AP54, AP57, Bet50, BHB^{+}51, Bri54b, Bri54a, Hae51, Ras58]. Alpha-Disintegration [AP54, AP57]. Alpha-Particle [Hae51, Ras58]. Alpha-Particles [Bet50, BHB^{+}51]. Alpha-Radioactivity [Ras58]. Aluminum [BT58a, BS58c, BS58d]. Aluminum-Rich [BT58a]. Alums [Eis52]. American [Bet55, Ano57b]. Amplitude [KM58]. Analog [KW57]. Analysis [CDLR55, Len59, SMLK55, Sto53]. Angular [All59a, BR53, BB52, Tem58, Van51]. Anisotropic [Chi53, Hel59]. Anisotropy [Kay53, SM53]. Annihilation [Fer56]. Announcement [Ano57a]. Anomalous [Ble53a]. Anthracene [Cru58, Rob58]. Antiferromagnet [Fin53]. Antiferromagnetic [Gor53b, Gor53d, MM53, Sno53b]. Antiferromagnetism [Gor53b, Kub53, Née53a, Sma53]. Antiproton [Seg58]. Apparatus [dKH53]. Application [Kub53, Pos56, Sut59]. Applications [Hog53, Sei55]. Approaches [Mei59]. Approximate [HL51]. Approximation [NB57, RS57]. April [Bet55]. Artificially [Mit50]. Aspects [DS54, Lev59a, Ros59b, Woo59]. Assessment [Whe57]. Associated [BBR52]. Associations [Amb58]. Astronomical [Cle57]. Atmosphere [EF56]. Atomic [CDLR55, DC53, DHP54, Duc57, GGMT54, Gol53, Har58, Meg58, RTG58, Wex58, Zin55]. Atoms [All58, All59b, Hen58, Lav58, Lon58]. Attempt [BT58a]. Attenuation [KM58, LT53]. Attraction [McV58]. Axially [RS57]. Axis [Ing57].

B [Kih53a, Kih55]. Balance [Wat55a]. Band [Her58, Lax58, Par59, Sla53]. Bands [Kas59]. Barium [vH50]. Barkhausen [Gor53a]. Basis [Coh53, Sch59b]. Beams [Wex58]. Benzene [Cox58, Cru58]. Beryllium [Hae51]. Beta [All59a, FT50, Fei51, Gol59, KP52, Kin54, Kon55, Lid57, MMN51, Mic57, Nor51, Wu59]. Beta-Decay [All59a, FT50, Fei51, Kon55, MMN51, Nor51]. Beta-Disintegration [Kin54, Lid57]. Beta-Ray [KP52]. Between [YHA53, Kas59, Mic57, Sto59, Wei59b]. Biochemical [Ano59d, Rob59]. Biological [Ano59a, Ano59k, FM59, Kau59, Ric59b, Sch59d, Sut59]. Biologically [Dot59a]. Biology [Ano59b, Gli59, Sch59b]. Biophysical [ABB^{+}59]. Biosynthesis [Kor59, Mei59]. Bismuthide [Ada53]. Blood [Wau59a]. Bond [Org59]. Bonding [Pep58]. Bonds [Meg58]. Bone [Gli59]. Born [Koh54b, Koh54a]. Both [McV58]. Bound [JR58, JR59]. Bremsstrahlung [JR59, JR58, KM59]. Bright [Pot58]. Broadening [yCT57]. Bromine [Smo53b]. Bromine-Graphite [Smo53b]. Brownian [Cox52]. Bureau [dKH53].

C [SMLK55]. CaF [WW58]. Calorimetry [AN59]. Capture [BR58, Che56, Gol59, MB54, Pri59, RF55, BR59]. Carbohydrates [Onc59]. Carbon [Hae51]. Carrier [Tau57]. Cell [Ben59a, LW57, LW59]. Cells [Ben59b, Puc59, Tob59, Wei59b]. Cellular [Ano59b, Blo59, Wei59a, Woo59]. Centers [Sei54]. Central [Bul59]. Certain [Sed58]. Chains [Fes58]. Change [Cox50]. Changing [All58, All59b]. Chapel [Ber57]. Characteristic [MH58]. Characteristics [Ano59i]. Characterization [Dot59b, Onc59]. Charge [All58, All59b, KW57, Rav58, Wat55b]. Charge-Changing [All58, All59b]. Chemical [Ano59i, Ano59k, Kau59, Onc59, Ver59]. Chloroplasts [Hod59b]. Chromium [Eis52]. Chromosomes [Ben59a]. Cilia [Ino59]. Cl [Gor53d]. Classical [DeW57a]. Cloud [Don58, Saf58]. Clouds [Dav58, Kah58a, Kap58a, Roj58]. cm [Dav58, Ker58]. CO [Hag51, SB58, WD58]. Coagulation [Wau59a]. Cobalt [Gui53]. Coding [Ell59, Lev59a, Sch59d]. Coefficient [KR52]. Coefficients [BBR52, DE52, Kih53b, Kih55]. Coercive [Gei53a, Mei53]. Coherence [KB58]. Coherent [LS58]. Coils [LK53b]. Collagen [Sch59a]. Collective [Mot57, Pin56, Woh53]. Collision [Kah58a, Sea58b]. Collisions [All58, All59b, Lew52, Sav58b]. Color [Sei54]. Comment [Sle60, Wil60]. Comparative [Che56, FT50, MB54, Var57, Var59]. Comparison [AN59, Zan58]. Compounds [Rat53]. Compressibility [Bru58]. Compressional [Bor53]. Compton [Cri58, Cur58]. Computer [Utt59]. Concentrated [Zim59]. Concentrations [Tau57]. Concerning [Meg58]. Conclusions [Ano58b]. Condensation [Sch58a]. Conditional [Utt59]. Conduction [Kas59]. Conference [Ano59f, Mar56a, Sla58a, Ber57]. Configuration [Gol57, Ric59b]. Configurations [BT58a, Dot59a]. Congress [Con57, Hof58]. Conjugation [Wat55b]. Connective [Ano59c]. Conservation [Hil51]. Constants [AN59, CDLR55, DC53]. Containing [LP58, WvdH53]. Contracting [Mes58]. Contraction [Mor59]. Contributions [RSB53]. Control [Ano59k]. Controlled [Pos56]. Convection [AB58]. Convergence [Koh54b, Koh54a]. CoO [Fin53, NM58]. Cooling [Sch58b]. Coordinates [RRS54]. Copper [Ble53a, Bra51]. Correlation [All59a, BR53, Cam58]. Correspondence [Kur57]. Cosmic [BBC^{+}52, Mor57]. Cosmic-Ray [BBC^{+}52]. Cosmical [Bos58]. Cosmogonical [Haa50]. Cosmological [BW57a]. Coulomb [BHB^{+}51, Frö55]. Coupling [Van51, Van53, YHA53]. Cr [Hen53, WD58]. Crab [Mün58b, Shk58]. Criterion [Coh53]. Critical [EM58, Low58]. Cross [Ada50, Bar52, BB52, BS58e, Elt58, Hes58, KM59]. Cross-Section [Hes58, KM59]. CrSb [Sno53b]. Crystal [Bel54, Cox58, Ell53, Gor53b, Kay53, Lav58, Sla58a]. Crystals [Bac58, Cur58, GAH53, Gor53d, GW58, Le 58, LT53, LP58, Lon58, MH58, Sei51, Sei54, Sla58b]. Cu [Gor53d, Hen53, WD58]. Current [Sla58a, Tau57]. Curved [DeW57a]. Curves [Sto53]. Cygnus [Min58]. Cylindrical [WW57]. Cytoplasm [Sjö59].

Dark [Fes58]. Data [CDLR55, Hes58, KM59, WS53]. Dayton [SMLK55]. Debye [AN59]. Decay [All59a, Dal59, FT50, Fei51, Gol59, Kon55, MMN51, Mic57, Nor51]. Decays [Wu59]. December [Hof58]. Deformation [Bri50]. Demagnetization [dKH53]. Density [Fan57, Gla58, GW58, Hof58, Sea58a, vdH58]. Deposition [Wex58]. Description [Fan57]. Determination [All59a, Bac58, GGMT54, KP52, WD58]. Determinations [Coc58, KH58, RTG58]. Determine [BT58a]. Determined [AN59, Fer58]. Deuterium [Mar51]. Deuteron [Ma53]. Development [Ano57b]. Developments [Roo51]. Diagram [Che56, MB54]. Diamagnetism [Mye52a, Mye52b]. Diatomic [MMMW58, Var57, Var59]. Dielectric [MWV57]. Differences [DHP54, Duc57]. Diffraction [Bac58, Coc58, HC53, LP58, SW53, WS53]. Diffuse [Jac58, Pot58]. Dimensional [Ken59]. Dioxide [Gra59]. Dirac [Goo55]. Direct [Bro59, Wat58]. Direct-Interaction [Wat58]. Direction [dV58]. Discharges [Kih52, Kih53a]. Discussion [Abr53, Ano58a, BB53, Ble53b, CE53, FW53, Gei53b, GGHG53, GHS^{+}53, GA53, GG53, Gor53c, GZ53, GS53, GSB^{+}53, KG53, LK53a, ME53, OWL53, Rad53a, Rad53b, RK53, RBFF53, RHWR53, Smo53c, Sno53a]. Disintegration [AP54, AP57, Kin54, Lid57, VW54, VW57]. Disk [Ker58]. Dispersion [Le 58, WGvdB53]. Dissipation [BS58b, Pet58]. Distortion [AB58]. Distribution [AH50b, AH50c, BB52, Coc58, EJ58, GW58, Kih53a, vdH58]. Distributions [Cam58, Gla58, Hof58, LS58, Rav58, WW58]. Divergence [Des57]. Domain [DM53, GAH53, vH50]. Driven [BC59]. Due [yCT57, Ver59, Whi55]. Dynamical [DeW57a, Sav58a]. Dynamics [Jac58, LW57, Par58, Sav58b, Wei59a, Zan58, LW59]. Dynamo [Els56]. Dysprosium [LSBE53].

Earth [Els50, Ing55b, Ing57]. Earths [Ell53]. Effect [Cri58, Hog53, PR53, WvdH53]. Effects [Ano59a, BM53, Dam53, EJ58, Fin53, Ope53, Sav58a, Tob59].

Eigenfunctions [Mel56]. Eigensolutions [HL51]. Einstein [Gup57, Opp56, WW57]. Elastic [AN59, Le 58]. Electric [KW57]. Electrical [Kih52, Kih53a, Le 58, LSBE53, Ros59b]. Electricity [Lan57a]. Electrolytes [Ver59]. Electromagnetic [ABH^{+}56, ABH^{+}58, HBY58, Sch58c, YLR57]. Electromagnetism [BCK57]. Electron [All59a, Bar51, BT58a, BS58e, Che56, Coc58, Fol58, Gab56, GW58, Hag51, Hof56, MB54, Mar59, Mar56a, Mar56b, Pag59, Rav58, RF55, SO55, Sim56, Tol56, WW58, Woh53]. Electron-Capture [RF55]. Electron-Neutrino [All59a]. Electron-Vibration [Bar51]. Electronic [Her58, Lax58, Par59, RSB53]. Electrons [BCK57, JR58, KP52, Mas56, WD58, JR59]. Elementary [FV58, Gui53, Hei57, Sha56]. Elements [Ada50, AH50c, AP54, AP57, BBFH57, Eis54, SW53, SU56, AH50b]. Elongate [Sch59a]. emf [Tau57]. Emission [Shk58]. Emphasis [Ros59b]. Empirical [Meg58]. Emulsion [Bei52]. End [KP52]. End-Point [KP52]. Energies [All58, All59b, AP54, AP57, Bri54b, Bri54a, KP52, Kin54, Lid57, Nie51, RF55, VW54, VW57]. Energy [AL52, AL55, AH50a, Ano59d, Bet50, BFGS58, Bro53, Cal59a, Cam58, EK54, EB57, GC59, Haa50, Hae51, Her58, Hes58, HLMF50, Ing53, Ing55a, JB50, KP52, Leh59, Mar56b, Pin56, Sal57, Set58, Var57, Var59, Zin55]. Enrico [Bet55, And55]. Entropy [Bri50]. Enzyme [Alb59, Hog59, Neu59]. Enzymes [Pol59]. Equations [EF56, Inf57, McV58]. Equivalence [Dic57a, Dic57b]. Erbium [LSBE53]. Errata [ABH^{+}58, All59b, BW61, BR59, JR59, VV59, Var59]. Erratum [AH50b, Bri54b, BS58c, Ing55a, Koh54a, LW59, SH50, Web60]. Error [Utt59]. Estimates [Sha58]. Estimation [ACP53]. Ethyl [RSB53]. Even [Nor51]. Everett [Whe57]. Evidence [BFGS58]. Exact [McV58]. Examples [Sed58]. Exchange [AW53b, Kit53, Ste53, Van53, YHA53, ZH53]. Exchange-Introduction [Kit53]. Excitation [ABH^{+}56, ABH^{+}58]. Exciton [Kas59]. Expansion [Chi53, Sno53b]. Expansions [Koh54b, Koh54a]. Experiment [Tol56]. Experimental [All58, BH56, BFGS58, Gor53a, Lax58, Mei53, Nus56, Woh53, All59b]. Experiments [All59a, And55, Gab56, Mar56b, Sim56, Sle60, Wil59a, Wil60, Wu59]. Eye [Har59].

Factorization [IH51]. Far [BBC^{+}52]. Faraday [Hog53]. Fast [BRTW52, Bar52]. Fe [Hen53, WD58]. Features [Zir59]. Fermi [Bet55, And55, AA55, Dal59, FT50, Kon55, Seg55, Sei55, Zin55]. Ferric [WvdH53]. Ferrite [GAH53, HC53]. Ferrites [Bic53, PS53, Rad53c, WGvdB53, WvdH53]. Ferroelectricity [vH50]. Ferromagnetic [CH53, Dam53, Hog53, MWV57, PR53, Van53, Wei53]. Ferromagnetism [KA53, Née53a, NM53a, NM53b, Sla53, Woh53]. Ferrous [WvdH53]. Feynman [Mis57]. Fiber [Sch59c]. Fibrous [Hod59a]. Field [AB58, Des57, Ell53, JR58, JR59, Kih53a, Mes58, RS57, Ser58, Sma53]. Fields [BCK57, BW57b, Hei57, Hen53, LK53b, McV58, BW61]. Filaments [Fes58, Mün58b]. Films [CH53]. Finding [Mel56]. Fine [Ben59a, FM59, Hod59b, Len59, Mei53, Née53b, Sha58, Sjö59, Wei53]. Fine-Structure [Len59]. First [RS57]. Fission [Cam58]. Fluids [MBL51]. Fluorescence [RF55]. Fluorides [SM53]. Fluorite [Cri58]. Flux [BRTW52]. Force [Gei53a, Mei53]. Forces [Sto59]. Foreign [yCT57]. Formation [Kap58a, OS58, Sav58a]. Formula [Gre58]. Formulas [Bro53, KM59]. Formulation [Eve57, Whe57]. Foundations [Bro59, Haa55, Lan56]. Fourier [WS53]. Franck [Pri52]. Free [Cal59b, JR58, JR59, Mar56a]. Frequencies [HL53, Hog53, Mas53]. Frequency [WGvdB53]. Fronts [Gol58, Kah58b]. Function [Ano59g]. Functions [BHB^{+}51, Frö55, Har58, Var57, Var59]. Fund [Ano57b]. Fusion [Pos56].

Gadolinium [Kip53, LSBE53]. Galactic [Fes58, Ker58, PS58, Ser58]. Galaxies [dV58]. Galaxy [Sav58b]. Gamma [DE52, FH59]. Gamma-Ray [DE52]. Gap [BFGS58]. Gas [Kah58a, Kap58a, SO55, Sav58a, Sav58b, Sed58, Zan58, vdH58]. Gas-Dynamical [Sav58a]. Gases [yCT57, KR52, Kih52, Kih53a, Kih53b, Kih55]. Gatlinburg [Ano59f]. Gd [Hen53]. General [Ano58a, Rob59, Zir59, Bon57, Des57, Inf57, Lan57a, Mis57]. Generalized [Bri50, Kur57]. Generation [Tau57]. Genetic [Len59]. Genetics [Ano59e, Lev59b]. Geomagnetism [Els50]. Geons [Ern57, PW57]. Gibbs [Cox50]. Glancing [FB51]. Globular [Ken59]. Globules [Roj58]. Gradients [LK53b]. Graphite [Smo53b]. Gravitation [Dic57a, Gup57, Kur57]. Gravitational [BS58a, BW57b, BW61, RS57, Sch53, WW57]. Gravity [BCK57]. Gross [Par58]. Group [Bel54, SM53]. Groups [Gol57]. Gyromagnetic [DB53, Eps56].

H [Gor53d]. Half [Che56, FT50, MB54]. Half-Lives [Che56, FT50, MB54]. Halide [Sei51, Sei54]. Hall [PR53]. Halo [PS58]. Hamilton [Hil51]. Having [MH58]. Heat [BM53, RSB53]. Heavy [AW53a, AP54, AP57]. Held [Ano59f, Bet55, Hof58]. Helium [All58, All59b, DS54]. Helium-Some [DS54]. Helmholtz [LK53b]. Hexagonal [Rat53]. HI [Ker58]. High [Hes58, Kri53, Mas53, Pos56, Sle60, WGvdB53, Wil59a, Wil60]. High-Energy [Hes58]. High-Temperature [Sle60, Wil59a, Wil60]. Highly [Kah58a]. Hill [Ber57]. Homeostasis [Wau59a]. Honor [Bet55]. Hormone [Ste59]. Hydrogen [All58, All59b, Bac58, LP58, Mar51, Org59, Pep58]. Hydromagnetic [Bos58, Els56]. Hyperfine [EJ58, Whi55]. Hypotheses [Sed58].

II [Ano58a, Duc57, EB57, Lid57, Nor51, Sei54, VW57, Wat55b]. III [Ano58a, HLMF50, Wat55c]. Illustrated [Hod59b]. Imitation [Utt59]. Impact [Hag51]. Implications [Wat58]. Important [Dot59a]. Impulse [Kat59b]. Impulses [Bul59]. Inactivation [Pol59]. Incidence [FB51]. Induced [BLR53, Hog59]. Inelastic [Sea58b]. Information [Ell59, Har59]. Infrared [Sut59]. Initiation [Bul59]. Instability [BS58a]. Institute [Ano57b]. Integral [MB56]. Integration [Har59]. Intensity [MH58]. Interacting [BCK57]. Interaction [All59a, BW57b, Bro59, Dal59, Fol58, Mar59, Sch59a, Sla58b, Wat58, BW61]. Interactions [Ano59f, Bar51, BH56, Dic57b, GM59, Mar56a, Mic57, Ste53, Wau59b, Wei59b, ZH53]. Interference [Gab56, Sim56]. Interferometer [SMLK55]. Interferometry [KB58]. Interior [Els50]. Internal [LS59, Mün58a, Wil58b]. International [Con57, Sla58a, Hof58]. Interpretation [BBC^{+}52, Elt58, FT50, JB50, Ma53]. Interstellar [Dav58, Don58, Kap58a, Par58, Sch58b, Sch58a, vdH58]. Introduction [BJ53, Kit53, Smo53a, Wic55]. Introductory [DeW57b, Yuk57]. Invariance [Wig57]. Inversion [WS53]. Inversions [Wat55b]. Investigation [Her58, Van57, sW50]. Investigations [Lax58]. Invited [Ano59f]. Ionization [Gol58, Hag51, Kah58b]. Ionized [Kah58a]. Ions [ABH^{+}56, All58, All59b, Kih53a, KW53, Mye52a, Mye52b, WvdH53, YHA53, ABH^{+}58]. Iron [BM53, DM53, Gor53a, Gui53, Low58, Rat53, SM53, YHA53]. Iron-Group [SM53]. Irreversible [BC59, Cox50, Cox52, Dom54]. Ising [NM53a, NM53b]. Isomerism [GH52, SH49, SH50]. Isotope [Bre58, BK58]. Isotopes [Hae51, HPS53, SHS58]. IV [AL52].

James [Pri52]. January [Mac50]. John [Ano51].

K-Capture [RF55]. K-Electron [Che56]. kev [All58, All59b]. Kinematics [Mün58b, Wil58b]. Kinetic [All58, All59b, Cam58]. Knowledge [Yan57].

L [RF55]. L-Fluorescence [RF55]. Laboratory [Bos58]. Lamellar [FM59, Hod59b, Kas59]. Large [dV58]. Large-Scale [dV58]. Lattice [Cru58, Jac58, LW57, LW59]. Lattices [Bel54, MMMW58]. Laws [Wat55a, Wat55b, Wat55c]. Layers [Sjö59]. Learning [Utt59]. Least [Coh53, DC53]. Least-Squares [DC53]. Lectures [Foc57]. Level [Ma53, Nus56]. Levels [AL52, AL55, AH50a, EK54, EB57, Hae51, HLMF50, Ing53, Ing55a]. LiF [WW58]. Life [Sch59b]. Light [AL52, AL55, EK54, EB57, Hel59, HLMF50, Ing53, Ing55a]. Lighter [BHB^{+}51]. Line [Bre57]. Linear [Ern57, MB56]. Linear-System [MB56]. Lines [yCT57, ML59, Van57]. Lipids [Onc59]. Liquid [DS54, HP53]. Liquids [BH56, KR52]. Lives [Che56, FT50, MB54]. Local [Sea58a]. Loop [Min58]. Losses [Mar56b, Pin56]. Low [Bic53, Eis52, Hen53, JB50, Mar56b, Set58, WB53]. Low-Energy [Mar56b, Set58].

Mach [FB51]. Macromolecular [Fer59, Zim59]. Macromolecules [Ano59i, Dot59a, Dot59b, Sch59a, Sto59]. Magnet [Ada53]. Magnetic [Bro53, Hen53, Kri53, LSBE53, LT53, Low58, McV58, Mes58, NM58, NH53, RRS54, Rad53c, SO55, Ser58, Smo53b, Sno53b, SM53, Whi55, WB53]. Magnetism [Ing55b]. Magnetite [WB53]. Magnetization [BM53, BLR53, Bro53, DB53, EJ58, Gui53, Hen53, Née53b, Rat53, Sto53, SCOT53]. Magneto [Ope53]. Magneto-Optical [Ope53]. Magnetogasodynamic [Kap58b]. Magnetohydrodynamic [BS58b]. Magnetohydrodynamics [McV58]. Magnetomechanical [Fin53]. Mammalian [Puc59]. Manganese [Ada53]. Masers [Web59, Web60]. Mass [Bon57, DHP54, Duc57, Gre58, RTG58]. Masses [GGMT54]. Material [Sch58a]. Materials [CH53, PR53]. Mathematical [Bas58, Hil51, Kih52, Kih53a]. Matrices [Goo55]. Matrix [Fan57, LT58]. Matter [AW53a, LS58, Sch58b]. Mean [RF55]. Measurement [BRTW52, RSB53]. Measurements [BBC^{+}52, FH59, Seg58, Sha58, SCOT53]. Measuring [Bar52, DB53]. Mechanics [Eve57, Fan57, FV58, Haa55, Van57]. Mechanism [Amb58, Ino59, Sed58]. Mechanisms [Alb59, Gui53, Har59, Kat59a, Mor59, OS58]. Media [Van53]. Medium [BS58a, Don58, Eps56, Par58]. Meeting [Bet55]. Membrane [Ben59a, Ber59]. Membranous [Sjö59]. Memorial [Bet55]. Mesodynamics [Low57]. Meson [And55, Gol59, Lew52, Mar51, MS51, Mil55, Wic55, Yuk57]. Mesonic [Hen58]. Metabolism [Con59]. Metal [Kip53]. Metals [BT58a, Chi53, LSBE53]. Method [Cox50, DB53, IH51, Kub53, LW57, LW59]. Methods [Bar52]. MeV [FB56]. Microstructure [Ser58]. Microwave [GGMT54, HL53, Hog53, Kip53, LS59, RTG58]. Miller [SMLK55]. Mineralized [Gli59]. Mirror [KH58]. Mishima [NH53]. Mitosis [Ino59]. Mixed [Gui53]. Mn [SB58]. Mo [Kri53]. Model [Bro59, Don58, Hae51, NM53a, NM53b, Wat58]. Models [BW57a, Jen57, Kih53b, Kih55, Van53]. Modes [BS58c, BS58d, Mar59]. Molecular [Ano59g, Gli59, Kas59, Ric59b, Ric59c, Roo51, Sch59b, Sch59c, Sma53]. Molecules [Kih53b, Kih55, Lon58, Nie51, Van51, Var57, Whi55, YHA53, Var59]. Moment [Sno53b]. Moments [BS56, Gol53, Mac50]. Momentum [Van51]. Monoenergetic [FB56]. Motility [Ino59]. Motion [Cox52, EF56, GAH53, Inf57, Mot57, Sch53, Sed58]. Motions [Mün58a]. Mu [Gol59]. Multicellular [Tob59]. Multiple [Lax51, Lew52, Mil55, OS58]. Muon [Pri59]. Muscle [Ano59c, Ben59b, Hod59a, Mor59].

N [Hag51]. NaCl [WW58]. Naphthalene [Cru58, Rob58]. Narrowing [AW53b]. National [dKH53]. Natural [Bri54b, Bri54a]. Nature [Kat59b, PS58, Shk58]. Nb [Kri53]. Nd [KW53]. near [RSB53]. Nearly [FB51]. Nebula [Mün58a, Mün58b, Shk58]. Nebulae [Fes58, Pot58, Sea58a, Wil58b, Zan58]. Nebulosities [Min58]. Negative [Bon57]. Neodymium [RSB53]. Nerve [Ano59h, Bul59, Con59, Kat59b, Sch59c]. Nervous [Ros59b]. Neurons [Bul59]. Neutral [LS58]. Neutrino [All59a, KW57]. Neutrinos [BW57b, BW61]. Neutron [Ada50, Bac58, BRTW52, Bar52, BS58c, BS58d, Fol58, FB56, HC53, HBY58, LP58, Pep58, Sch58c, Seg55, Set58, SW53, SB58]. Neutron-Electron [Fol58]. Neutrons [Fer58, KW53, Low58, NM58]. NH [HP53]. Ni [WD58]. Nickel [Gor53a, HC53, Mas53, WGvdB53]. Nickel-Iron [Gor53a]. NO [Hag51]. Noise [BC59, Gor53a]. Nonequilibrium [Tau57]. Nonlinear [BC59]. Nonuniform [BS58a]. Normal [BS58d, Mar59, BS58c]. Note [DeW57b, Ste53, WGvdB53]. Notebooks [AA55]. Nuclear [AH50a, ABH^{+}56, ABH^{+}58, Bei52, BR53, BLR53, BS56, Bru58, EJ58, Elt58, FH59, Fer58, Gla58, GH52, Gre58, Hen58, Hof56, Hof58, Jen57, Kli52, KH58, Lan57b, LT58, LS58, Mac50, MMN51, Nor51, Ras58, Rav58, RSB53, SH49, Seg58, Set58, Sha58, VW54, VW57, Wat58, Wei57, Wil58a, SH50]. Nuclei [AL52, AL55, BHB^{+}51, BK58, Bro59, EK54, EB57, HLMF50, Ing53, KH58, MMN51, Mit50, Nor51, Tem58, Ing55a]. Nucleic [Ano59e, Kor59, Lev59b, Pol59, Ric59c, Wil59b]. Nucleoli [Ben59a]. Nucleon [BG58, Hes58, Lew52]. Nucleon-Nucleon [Hes58, Lew52]. Nucleons [YLR57]. Nucleus [Ben59a, Mot57]. Nuclides [Che56]. Number [Gol57, Har58, SO55, Wat55b, WD58]. Numerical [Frö55].

O [Gor53d, Hag51]. Observables [NB57]. Observations [Gor53d, Ker58, SMLK55]. Ocean [EF56]. October [Ano59f]. Odd [Che56, MMN51]. Old [Mic57]. one [Dav58]. Opening [Whi58]. Operator [Fan57]. Optical [Bro59, Le 58, Ope53]. Optically [BLR53]. Orbital [BR58, BR59, RF55, Roo51]. Organic [GC59, YHA53]. Organisms [Ros59a, Tob59]. Organization [Ano59g, Sch59c, Sjö59]. Orientation [Sno53b]. Origin [AH50b, AH50c, Amb58]. Origins [Mor57]. Orion [Mün58a]. Other [Gup57]. Outbursts [Sed58]. Oxide [Bra51, Rat53]. Oxides [Gui53].

Pair [JR58, JR59]. Papers [Ano59f, Ano59a, Ano59b, Ano59c, Ano59d, Ano59e, Ano59g, Ano59h, Ano59i, Ano59k, Ano59j]. Paramagnetic [AW53b, HL53, HP53, KW53, Ope53]. Paramagnetism [Ble53a]. Parity [Wu59]. Part [Ano58a, Ano58b, Wat55a, Wat55b, Wat55c]. Particle [Dal59, Hae51, Nus56, Ras58]. Particles [AW53a, Bet50, BHB^{+}51, Bri54a, FV58, Hei57, Mei53, Mic57, Pol59, Sha56, Yan57, Bri54b]. Particular [Gli59, Sch59a]. Parts [Ano58a]. Passage [AW53a]. Pattern [DM53, Utt59]. Patterns [Rob59]. Performance [Ros59a]. Permalloy [Boz53, Kay53]. Permanent [Ada53]. Phage [Pol59]. Phase [vH50]. Phenomena [Bos58, PS53, Wig57]. Phosphors [GC59, HL53]. Photoionization [BS58e]. Photosynthesis [Cal59a]. Photosynthetic [Cal59b]. Physical [Ano59i, Bet55, Eis52, NH53, Sch59b, Wat55a, Wat55b, Wat55c]. Physicochemical [Dot59b]. Physics [ACP53, Ano57b, Con57, Hil51, Pos56, Seg55, Sla58a, Wei57]. Pi [Gol59, Low57]. Pi-Meson [Gol59]. Pion [LS58]. Plane [Bor53]. Planetary [Sea58a, Wil58b, Zan58]. Plasma [Sle60, Wil59a, Wil60]. Plasmas [ML59, Pos56]. Plastic [Bri50]. Point [KP52]. Polarization [FH59, Le 58, Pag59, Tol56]. Polyelectrolytes [Ric59a]. Polymers [Ric59b]. Populations [Tob59]. Positions [Bac58]. Positive [Kih53a]. Positron [BH56, Fer56, Pag59]. Possible [Bos58, OS58]. Potential [Var57, Var59]. Powdered [Ada53]. Powders [Née53b, Wei53]. Powerful [LK53b]. Precipitation [Gei53a]. Prediction [Wat55c]. Preface [BT58b]. Presence [yCT57]. Present [McV58, Yan57]. Pressure [Bor53, Zan58]. Principle [Cas57, Dic57a, Dic57b, Inf57, Hil51]. Principles [DeW57a]. Probability [Utt59]. Problem [Amb58, Boz53, DS54, Des57]. Problems [Ano58a, Cas57, Haa50, Meg58, RRS54, Sla58a, Sut59, Zan58]. Process [KA53]. Processes [BC59, Con59, Cox52, Dal59, Sch59b, Sea58b]. Processing [Ano59j]. Production [JR58, JR59, LS58, Lew52, LK53b, Mil55]. Professor [AA55]. Program [ABB^{+}59]. Progress [BLR53]. Projected [Roj58]. Propagation [Eps56, Gol58]. Properties [BBR52, Eis52, Goo55, Gra59, LSBE53, SO55, Sch59a, Sei51, Sha56]. Protein [Lev59a, Neu59, Sch59a]. Proteins [Ano59e, Hod59a, Ken59, Mei59, Onc59, Wau59b]. Proton [Gla58, HBY58, JB50]. Proton-Proton [JB50]. Protons [Bet50, BHB^{+}51]. Protoplanetary [Saf58]. Protostar [Mes58].

Q [Gol57, Wat55b]. Q-Number [Gol57, Wat55b]. Quantifiable [Ros59b]. Quantitative [Puc59]. Quantization [BCK57, Mis57]. Quantum [DeW57a, Des57, Eve57, Fan57, Hei57, Mar56a, Whe57, Wig57]. Question [KW57].

R [LT58]. R-Matrix [LT58]. Racah [BBR52]. Radiation [Ano59a, Bor53, Pol59, Zan58]. Radiations [BR53, Tob59]. Radicals [Cal59b]. Radii [Fer58, Hen58, Ras58, Seg58, Set58]. Radioactive [Mit50]. Radioactivity [Ras58]. Radiobiological [Zir59]. Radiobiology [Woo59]. Raised [Ano58a]. Range [Bet50, FB56, KP52]. Range-Energy [Bet50, KP52]. Rare [Ell53]. Ratios [DB53, RF55, RTG58]. Ray [BBC^{+}52, Coc58, Cri58, DE52, KP52, Lav58, MH58, Par59, Sha58, WD58, sW50]. Rays [Cur58, FH59, Jac58, Mor57]. Reaction [BB52]. Reactions [BHB^{+}51, Lan57b, LT58, Mar51]. Reality [WW57]. Reception [Ano59j, Cal59a]. Receptor [Bul59, Har59]. Reciprocity [Cas57]. Recognition [Utt59]. Recovery [Con59]. Rectifier [Bra51]. Reference [Gli59, Sch59a]. Reflection [FB51, MH58]. Region [Nus56]. Regulation [Ste59]. Related [KM59, Min58, Ano59a, Ano59b, Ano59c, Ano59d, Ano59e, Ano59g, Ano59h, Ano59i, Ano59k, Ano59j]. Relation [Bet50, Bro53, Bru58, Kas59]. Relations [KP52, MB56]. Relative [AH50b, AH50c, Eve57, Whe57]. Relativistic [FV58, Wig57]. Relativity [Bar57, Bon57, Des57, Foc57, Inf57, Lan57a, Mis57]. Relaxation [Dam53, KA53, Mas53, Ver59]. Remark [BW57a]. Remarks [MMMW58, Yuk57]. Renormalization [MS51]. Replication [Ano59e, Wil59b]. Reply [Wil60]. Report [BLR53]. Requirements [KB58]. Research [Pos56]. Resonance [AW53b, Dam53, Gor53b, HL53, HP53, Kip53, MM53, Ope53, PS53, RRS54]. Respiratory [Leh59]. Respiratory-Energy [Leh59]. Responses [Blo59, Ros59b]. Results [All58, All59b, BR58, BR59, Née53a]. Retrodiction [Wat55c]. Review [DeW57a, Sle60, Wil59a, Wil60]. Rheology [Fer59]. Rhombic [Gor53b]. Rich [BT58a]. Richter [WvdH53]. Rims [Pot58]. Role [Lev59b]. Rosen [WW57]. Rotating [RRS54, Sle60, Wil59a, Wil60]. Rotating-Plasma [Sle60, Wil59a, Wil60]. Rotation [BS58a, DB53, Ing57, LS59, Nie51, Whi55, dV58]. Rotation-by-Magnetization [DB53]. Rotational [Hae51, Mas53]. Rutile [Gra59].

Sargent [Che56, MB54]. Saturation [Bru58, Rat53]. Scale [dV58]. Scattering [BB52, Bre51, Cri58, Cur58, EM58, Elt58, Fer58, Gla58, Hel59, Hof56, JB50, Jac58, KW53, Lav58, Lax51, Low58, Mar56b, Mas56, NM58, Ras58, Rav58, Set58]. Schwarzschild [LW57, LW59]. Schwarzschild-Cell [LW57, LW59]. Science [ABB^{+}59]. Second [KR52]. Section [Hes58, KM59]. Sections [Ada50, Bar52, BB52, BS58e, Elt58]. Self [McV58]. Self-Attraction [McV58]. Semiconductors [MWV57, Tau57]. Semiempirical [Cam58]. Sensory [Ano59j, Har59, Ros59a, Ros59b]. Shape [Bre57]. Shapes [Tem58, sW50]. Shell [GH52, Jen57, Kli52, MMN51, Nor51, Nus56]. Shift [Bre58, BK58, yCT57]. Shifting [Ing57]. Shock [FB51, Kap58b]. Signal [Ano59j]. Silicon [BM53, DM53]. Silicon-Iron [DM53]. Silver [Sei51]. Simplification [Mel56]. Simulation [Bos58]. Single [LT53, Nus56, OS58]. Single-Particle [Nus56]. Singular [NB57]. Size [Bru58, Elt58, KH58, Sha58]. Sizes [Gre58, Hof58]. Slow [Bet50, Mas56]. Small [KM58]. Society [Bet55]. Solids [BH56, Fer56, Her58, KM58, Lax58, Par59, Pin56]. Solution [Dot59a]. Solutions [Hel59, HP53, McV58, Zim59]. Some [BBR52, Coc58, DS54, Dav58, Hen53, LSBE53, McV58, Meg58, Min58, Mit50, Née53a, Pep58, Ros59b, Wic55, Woo59]. Sound [MBL51, Ric55]. Sources [Sal57]. Space [Gol57, Wat55a, Wat55b]. Space-Time [Wat55a, Wat55b]. Spaces [DeW57a]. Spacings [Nus56]. Specific [RSB53]. Specificity [Ano59k, Kau59]. Spectra [Rad53c, SB58, sW50]. Spectral [yCT57, ML59]. Spectrometry [BS58d, SB58, BS58c]. Spectroscopic [DHP54, Duc57].

Spectroscopy [GGMT54, LS59, MWV57, Mit50, Par59, RTG58, Sut59]. Speculations [Sei51]. Spin [FV58, Kub53, VV58, VV59]. Spin-Wave [Kub53]. Spinors [BJ53]. Spontaneous [SCOT53]. Squares [Coh53, DC53]. Stability [Kah58b]. Standards [dKH53]. Stanford [Hof58]. Star [Fes58, OS58, Sav58a]. Stars [Amb58, BBFH57]. State [Eve57, Whe57]. States [BS58e, Fan57, SO55]. Static [Kih53a]. Statistical [ACP53, Cox50, Haa55, Mil55, Roj58, Van57]. Statistics [Sei55]. Status [GM59, Woh53]. Stellar [Amb58, Haa50, Roj58, Sal57, Sed58]. Stimuli [Ros59b]. Strange [Dal59]. Streaming [Hel59]. Stress [DM53, KM58]. Strong [Hen53]. Structure [ABH^{+}56, ABH^{+}58, Bas58, Ben59a, Ben59b, BG58, Cox58, EJ58, FM59, GH52, Gol58, Her58, Hod59b, Hof56, HBY58, Ing53, Ing55a, Ken59, Kli52, Lax58, Len59, ML59, MMN51, NH53, Neu59, Nor51, Par59, Ric59c, Rob58, Sch58c, Sha58, Sjö59, Wat58, Whi55, YLR57, dV58, vH50]. Structures [MH58]. Studies [BH56, BK58, Dav58, HC53, Hen53, LP58, Pep58, Puc59, SW53]. Study [ABB^{+}59, ABH^{+}56, Gor53a, Mei53, Roj58, Var57, Var59, Wau59a, WB53, ABH^{+}58]. Sulfate [RSB53]. Summaries [Ano58b]. Summarizing [Ano58a]. Summary [Ber57, Hes58, Smo53c]. Superconducting [Eis54]. Superconductivity [Bar51, Fey57]. Superconductors [BFGS58]. Superfluidity [Fey57]. Surface [Wil58a]. Survey [Van57]. Susceptibility [Kri53, Smo53b]. Suspensions [Hel59]. Sydney [Ker58]. Symmetric [RS57]. Symmetry [Mel56, Wat55a, Wat55b, Wat55c, Wat55a]. Symmetry-Adapted [Mel56]. Symposium [Bet55, Kit53]. Synaptic [Kat59a]. Synthesis [Ano59d, BBFH57, Hog59, Lev59a, Rob59]. Synthetic [Ric59b]. System [MB56, RS57, Ros59b]. Systematic [NB57]. Systems [Ano59k, BC59, Cal59b, FM59, Fer59, Hod59b, Kas59, Kau59].

T [Gor53b]. Table [AP54, AP57, Fei51, HPS53, Kin54, Lid57, Mac50, Sha56, SHS58]. Tables [Kli52]. Tate [Ano51]. Technique [Bei52]. Techniques [Fan57, FB56, SCOT53]. Temperature [Bic53, Kri53, MH58, Pos56, Sle60, Wil59a, Wil60, WB53]. Temperatures [Eis52, Hen53]. Tennessee [Ano59f]. Terms [Elt58]. Texture [Wei53]. Their [Wau59b, Bro53]. Theorems [Hil51, Wat55a]. Theoretical [BR58, Con57, Her58, Woh53, BR59]. Theories [BS56, Gup57, Ing55b, MS51, NB57, Wil58a]. Theory [AH50c, Bel54, BR53, Bre51, Bre58, Bro53, Cox52, DeW57a, Des57, Ell59, Ell53, EM58, Els56, Eps56, FT50, Fer56, Gab56, Hei57, Inf57, Kih52, Kih53a, Kon55, Kub53, Kur57, LT58, Mas56, Mil55, NM58, NM53a, NM53b, Pri59, Roo51, Sla53, Tol56, Wat55b, Whe57, Wic55, Woh53, Yuk57, AH50b, Foc57]. Thermal [Chi53, Cri58, Jac58, Lav58, Lon58, PW57, Sea58b, Sno53b]. Thermodynamics [BC59, Bri50, Dom54, Lan56]. Thermoelectricity [Dom54]. Thermoremanent [Née53b]. Thin [CH53]. Three [Foc57, Ken59, YHA53]. Three-Dimensional [Ken59]. Time [Cle57, Wat55a, Wat55b]. Tissue [Ano59c]. Tissues [Gli59]. Titanate [vH50]. Titanium [Gra59]. Topics [Bre51]. Toroidal [AB58, Ern57]. Torrence [Ano51]. Total [Kin54, Lid57]. Transducers [Sch59d]. Transfer [Ano59d, Cal59a, Cas57]. Transform [MB56]. Transformation [Bic53, Gol57, Leh59, WB53]. Transition [BT58a, RF55, SW53]. Transitions [Che56, MB54, vH50]. Transmission [Kat59a]. Transport [Ber59, GC59]. Traps [Mar59]. Treatment [Frö55, Sma53]. Trial [Utt59]. Trial-and-Error [Utt59]. Trivalent [KW53]. Tropocollagen [Sch59a]. Turbulence [Bas58, Kap58b, Saf58]. Twenty [Dav58]. Twenty-one-cm [Dav58]. Two [Kah58a, YHA53]. Type [MH58, WvdH53].

Ultrasonic [LT53, Ver59]. Underground [BBC^{+}52]. Uniaxial [Kay53]. Uniform [LK53b]. Universe [LW57, LW59]. University [Hof58]. upon [Ros59b]. Use [RRS54].

V [AL55, Kri53]. Valence [BS58e]. Valence-Electron [BS58e]. Validity [Don58]. Values [Fei51]. Vanadium [SB58]. Vapors [Ric55]. Variance [CDLR55]. Variation [DB53, Har58]. Variational [Kub53]. Variations [Sea58a]. Various [SW53]. Vectors [Van51]. Velocity [Ric55, vdH58, Kih53a]. Velocity-Distribution [Kih53a]. Vibration [Bar51, Nie51, SB58]. Vibration-Rotation [Nie51]. Vibrations [Cru58, Lon58, MMMW58]. VIII [Ano58b]. Virial [Kih53b, Kih55]. Virtual [Ma53]. Viscosity [KR52]. Vitro [Puc59].

W [Kri53]. Walls [GAH53]. Washington [Bet55]. Wave [Eps56, FV58, Frö55, Har58, Kub53]. Waves [Bor53, FB51, Kap58b, KM58, Lax51, Le 58, Sla58b, VV58, WW57, VV59]. Weak [Ano59f, Dic57b, GM59, Mic57]. Weiszäcker [Gre58]. without [Dic57a]. Work [Ker58, Wic55]. Wurtzite [MH58]. Wurtzite-Type [MH58].

X [Coc58, Cri58, Cur58, Jac58, Lav58, MH58, Par59, Pep58, Sha58, WD58]. X-Ray [Coc58, Cri58, Lav58, MH58, Par59, Pep58, Sha58, WD58]. X-Rays [Cur58, Jac58].

Yield [Cam58]. Yields [RF55].

Z [EK54, EB57]. Zinc [HC53, WGvdB53]. Zincblende [MH58]. Zincblende- [MH58].

References

Anderson:1955:PFN

[AA55] H. L. Anderson and Samuel K. Allison. From Professor Fermi’s notebooks. Reviews of Modern Physics, 27(3):273–275, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.27.273;   http://rmp.aps.org/abstract/
RMP/v27/i3/p273_1.

Allan:1958:DTF

[AB58] D. W. Allan and E. C. Bullard. Distortion of a toroidal field by convection. Reviews of Modern Physics, 30(3):1087–1088, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1087;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1087_1.

Abelson:1959:BSS

[ABB^{+}59] P. H. Abelson, J. W. Beams, R. H. Bolt, J. D. Ferry, I. Fuhr, D. E. Goldman, I. Gray, J. D. Hardy, H. K. Hartline, J. G. Kirkwood, H. Neurath, J. L. Oncley, M. D. Rosenberg, H. A. Sober, R. C. Williams, and Chairman F. O. Schmitt. Biophysical science — a study program. Reviews of Modern Physics, 31(1):1, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.31.1;   http://rmp.aps.
org/abstract/RMP/v31/i1/p1_1.

Alder:1956:SNS

[ABH^{+}56] K. Alder, A. Bohr, T. Huus, B. Mottelson, and A. Winther. Study of nuclear structure by electromagnetic excitation with accelerated ions. Reviews of Modern Physics, 28(4):432–542, October 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.28.432;   http://rmp.aps.org/abstract/

RMP/v28/i4/p432_1. See erratum [ABH^{+}58].

Alder:1958:ESN

[ABH^{+}58] K. Alder, A. Bohr, T. Huus, B. Mottelson, and A. Winther. Errata: Study of Nuclear Structure by Electromagnetic Excitation with Accelerated Ions. Reviews of Modern Physics, 30(2):353, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.353;   http://rmp.aps.org/abstract/
RMP/v30/i2/p353_1. See [ABH^{+}56].

Abragam:1953:D

[Abr53] A. Abragam. Discussion. Reviews of Modern Physics, 25 (1):268, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.268;   http://
rmp.aps.org/abstract/RMP/v25/i1/p268_1.

Annis:1953:SEP

[ACP53] M. Annis, W. Cheston, and H. Primakoff. On statistical estimation in physics. Reviews of Modern Physics, 25(4):818–830, October 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.818;   http://rmp.aps.org/abstract/
RMP/v25/i4/p818_1.

Adair:1950:NCS

[Ada50] Robert K. Adair. Neutron cross sections of the elements. Reviews of Modern Physics, 22(3):249–289, July 1950. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.22.249;
 http://rmp.aps.org/abstract/RMP/v22/i3/p249_1.

Adams:1953:NPM

[Ada53] Edmond Adams. A new permanent magnet from powdered manganese bismuthide. Reviews of Modern Physics, 25(1): 306–307, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.306;   http://
rmp.aps.org/abstract/RMP/v25/i1/p306_1.

Alburger:1950:NEL

[AH50a] D. E. Alburger and E. M. Hafner. Nuclear energy levels, Z — 11 to 20. Reviews of Modern Physics, 22(4):373–385, October 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.22.373;   http://rmp.aps.org/abstract/
RMP/v22/i4/p373_1.

Alpher:1950:ETO

[AH50b] Ralph A. Alpher and Robert C. Herman. Erratum: Theory of the Origin and Relative Abundance Distribution of the Elements. Reviews of Modern Physics, 22(4):406, October 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
22.406;   http://rmp.aps.org/abstract/RMP/v22/i4/p406_

1. See [AH50c].

Alpher:1950:TOR

[AH50c] Ralph A. Alpher and Robert C. Herman. Theory of the origin and relative abundance distribution of the elements. Reviews of Modern Physics, 22(2):153–212, April 1950. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.22.153;
 http://rmp.aps.org/abstract/RMP/v22/i2/p153_1. See erra-

tum [AH50b].

Ajzenberg:1952:ELL

[AL52] F. Ajzenberg and T. Lauritsen. Energy levels of light nuclei. IV. Reviews of Modern Physics, 24(4):321–402, October 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.24.321;   http://rmp.aps.org/abstract/
RMP/v24/i4/p321_1.

Ajzenberg:1955:ELL

[AL55] F. Ajzenberg and T. Lauritsen. Energy levels of light nuclei. V. Reviews of Modern Physics, 27(1):77–166, January 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
27.77;   http://rmp.aps.org/abstract/RMP/v27/i1/p77_1.

Alberty:1959:MEA

[Alb59] Robert A. Alberty. Mechanisms of enzyme action. Reviews of Modern Physics, 31(1):177–184, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.177;
 http://rmp.aps.org/abstract/RMP/v31/i1/p177_1.

Allison:1958:ERC

[All58] Samuel K. Allison. Experimental results on charge-changing collisions of hydrogen and helium atoms and ions at kinetic energies above 0.2 kev. Reviews of Modern Physics, 30(4): 1137–1168, October 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.1137;   http:
//rmp.aps.org/abstract/RMP/v30/i4/p1137_1. See erratum

[All59b].

Allen:1959:DBD

[All59a] James S. Allen. Determination of the beta-decay interaction from electron-neutrino angular correlation experiments. Reviews of Modern Physics, 31(3):791–796, July 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.791;
 http://rmp.aps.org/abstract/RMP/v31/i3/p791_1.

Allison:1959:EER

[All59b] Samuel K. Allison. Errata: Experimental Results on Charge-Changing Collisions of Hydrogen and Helium Atoms and Ions at Kinetic Energies above 0.2 kev. Reviews of Modern Physics, 31(3):839, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.839;   http://
rmp.aps.org/abstract/RMP/v31/i3/p839_2. See [All58].

Ambartsumian:1958:PMO

[Amb58] V. A. Ambartsumian. On the problem of the mechanism of the origin of stars in stellar associations. Reviews of Modern Physics, 30(3):944–948, July 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.944;
 http://rmp.aps.org/abstract/RMP/v30/i3/p944_1.

Alers:1959:CDD

[AN59] G. A. Alers and J. R. Neighbours. Comparison of the Debye θ determined from elastic constants and calorimetry. Reviews of Modern Physics, 31(3):675–680, July 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.675;
 http://rmp.aps.org/abstract/RMP/v31/i3/p675_1.

Anderson:1955:MEE

[And55] Herbert L. Anderson. Meson experiments with Enrico Fermi. Reviews of Modern Physics, 27(3):269–272, July 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.27.269;
 http://rmp.aps.org/abstract/RMP/v27/i3/p269_1.

Anonymous:1951:JTT

[Ano51] Anonymous. John Torrence Tate. Reviews of Modern Physics, 23(3):183–184, July 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.23.183;   http://
rmp.aps.org/abstract/RMP/v23/i3/p183_1.

Anonymous:1957:A

[Ano57a] Anonymous. Announcement. Reviews of Modern Physics, 29 (1):1, January 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.29.1;   http://rmp.aps.
org/abstract/RMP/v29/i1/p1_1.

Anonymous:1957:DFA

[Ano57b] Anonymous. Development fund of the American institute of physics. Reviews of Modern Physics, 29(2):254, April 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.254;   http://rmp.aps.org/abstract/
RMP/v29/i2/p254_1.

Anonymous:1958:PIG

[Ano58a] Anonymous. Part III: General summarizing discussion of problems raised in parts I and II. Reviews of Modern Physics, 30(3):994–1008, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.994;   http://
rmp.aps.org/abstract/RMP/v30/i3/p994_1.

Anonymous:1958:PVS

[Ano58b] Anonymous. Part VIII: Summaries and conclusions. Reviews of Modern Physics, 30(3):1095–1108, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1095;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1095_1.

Anonymous:1959:BER

[Ano59a] Anonymous. Biological effects of radiation (related papers: 32-35). Reviews of Modern Physics, 31(1):3, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.3;
http://rmp.aps.org/abstract/RMP/v31/i1/p3_1.

Anonymous:1959:CBR

[Ano59b] Anonymous. Cellular biology (related papers: 1-3, 47, 48). Reviews of Modern Physics, 31(1):2, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.2;
http://rmp.aps.org/abstract/RMP/v31/i1/p2_1.

Anonymous:1959:CTM

[Ano59c] Anonymous. Connective tissue and muscle (related papers: 41-46). Reviews of Modern Physics, 31(1):3, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.3;
http://rmp.aps.org/abstract/RMP/v31/i1/p3_3.

Anonymous:1959:ETB

[Ano59d] Anonymous. Energy transfer and biochemical synthesis (related papers: 16-22, 24, 25). Reviews of Modern Physics, 31 (1):2, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.31.2;   http://rmp.aps.
org/abstract/RMP/v31/i1/p2_3.

Anonymous:1959:GRP

[Ano59e] Anonymous. Genetics and replication of proteins and nucleic acids (related papers: 26-31). Reviews of Modern Physics, 31 (1):2–3, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.31.2;   http://rmp.aps.
org/abstract/RMP/v31/i1/p2_4.

Anonymous:1959:IPC

[Ano59f] Anonymous. Invited Papers from the Conference on Weak Interactions Held at Gatlinburg, Tennessee, October 27–29, 1958. Reviews of Modern Physics, 31(3):782, July 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.782;
 http://rmp.aps.org/abstract/RMP/v31/i3/p782_1.

Anonymous:1959:MOF

[Ano59g] Anonymous. Molecular organization and function (related papers: 36-40). Reviews of Modern Physics, 31(1):3, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.31.3;   http://rmp.aps.org/
abstract/RMP/v31/i1/p3_2.

Anonymous:1959:NRP

[Ano59h] Anonymous. Nerve (related papers: 49-51). Reviews of Modern Physics, 31(1):4, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.4;
http://rmp.aps.org/abstract/RMP/v31/i1/p4_1.

Anonymous:1959:PCC

[Ano59i] Anonymous. Physical and chemical characteristics of macromolecules (related papers: 4-15, 23). Reviews of Modern Physics, 31(1):2, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.31.2;   http://rmp.aps.
org/abstract/RMP/v31/i1/p2_2.

Anonymous:1959:SRS

[Ano59j] Anonymous. Sensory reception and signal processing (related papers: 52-58). Reviews of Modern Physics, 31(1):4, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.31.4;   http://rmp.aps.org/
abstract/RMP/v31/i1/p4_2.

Anonymous:1959:SCC

[Ano59k] Anonymous. Specificity in the chemical control of biological systems (related papers: 59-61). Reviews of Modern Physics, 31(1):4, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.31.4;   http://rmp.aps.
org/abstract/RMP/v31/i1/p4_3.

Asaro:1954:TAD

[AP54] Frank Asaro and I. Perlman. Table of alpha-disintegration energies of the heavy elements. Reviews of Modern Physics, 26 (4):456–462, October 1954. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.26.456;   http://
rmp.aps.org/abstract/RMP/v26/i4/p456_1.

Asaro:1957:TAD

[AP57] Frank Asaro and I. Perlman. Table of alpha-disintegration energies of the heavy elements. Reviews of Modern Physics, 29 (4):831–836, October 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.831;   http://
rmp.aps.org/abstract/RMP/v29/i4/p831_1.

Allison:1953:PHP

[AW53a] S. K. Allison and S. D. Warshaw. Passage of heavy particles through matter. Reviews of Modern Physics, 25(4):779–817, October 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.779;   http://rmp.aps.org/abstract/
RMP/v25/i4/p779_1.

Anderson:1953:ENP

[AW53b] P. W. Anderson and P. R. Weiss. Exchange narrowing in paramagnetic resonance. Reviews of Modern Physics, 25(1):269–276, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.269;   http://rmp.aps.org/abstract/
RMP/v25/i1/p269_1.

Bacon:1958:DHP

[Bac58] G. E. Bacon. Determination of hydrogen positions in crystals by neutron diffraction. Reviews of Modern Physics, 30 (1):94–95, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.94;   http://
rmp.aps.org/abstract/RMP/v30/i1/p94_1.

Bardeen:1951:EVI

[Bar51] J. Bardeen. Electron-vibration interactions and superconductivity. Reviews of Modern Physics, 23(3):261–270, July 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
23.261;   http://rmp.aps.org/abstract/RMP/v23/i3/p261_
1.

Barschall:1952:MMF

[Bar52] H. H. Barschall. Methods for measuring fast neutron cross sections. Reviews of Modern Physics, 24(3):120–132, July 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
24.120;   http://rmp.aps.org/abstract/RMP/v24/i3/p120_
1.

Bargmann:1957:R

[Bar57] V. Bargmann. Relativity. Reviews of Modern Physics, 29 (2):161–174, April 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.161;   http://
rmp.aps.org/abstract/RMP/v29/i2/p161_1.

Bass:1958:MST

[Bas58] J. Bass. On the mathematical structure of turbulence. Reviews of Modern Physics, 30(3):1084–1086, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1084;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1084_1.

Blatt:1952:ADS

[BB52] John M. Blatt and L. C. Biedenharn. The angular distribution of scattering and reaction cross sections. Reviews of Modern Physics, 24(4):258–272, October 1952. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.24.258;
 http://rmp.aps.org/abstract/RMP/v24/i4/p258_1.

Bleaney:1953:Da

[BB53] B. Bleaney and F. Bitter. Discussion. Reviews of Modern Physics, 25(1):177, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.177;   http:/
/rmp.aps.org/abstract/RMP/v25/i1/p177_1.

Barrett:1952:ICR

[BBC^{+}52] Paul H. Barrett, Lowell M. Bollinger, Giuseppe Cocconi, Yehuda Eisenberg, and Kenneth Greisen. Interpretation of cosmic-ray measurements far underground. Reviews of Modern Physics, 24(3):133–178, July 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.24.133;   http://
rmp.aps.org/abstract/RMP/v24/i3/p133_1.

Burbidge:1957:SES

[BBFH57] E. Margaret Burbidge, G. R. Burbidge, William A. Fowler, and F. Hoyle. Synthesis of the elements in stars. Reviews of Modern Physics, 29(4):547–650, October 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.547;
 http://rmp.aps.org/abstract/RMP/v29/i4/p547_1.

Biedenharn:1952:SPR

[BBR52] L. C. Biedenharn, J. M. Blatt, and M. E. Rose. Some properties of the Racah and associated coefficients. Reviews of Modern Physics, 24(4):249–257, October 1952. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.24.249;
 http://rmp.aps.org/abstract/RMP/v24/i4/p249_1.

Bernard:1959:ITN

[BC59] William Bernard and Herbert B. Callen. Irreversible thermodynamics of nonlinear processes and noise in driven systems. Reviews of Modern Physics, 31(4):1017–1044, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.31.1017;   http://rmp.aps.org/abstract/RMP/
v31/i4/p1017_1.

Belinfante:1957:QIF

[BCK57] F. J. Belinfante, D. I. Caplan, and W. L. Kennedy. Quantization of the interacting fields of electrons, electromagnetism, and gravity. Reviews of Modern Physics, 29(3):518–546, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
29.518;   http://rmp.aps.org/abstract/RMP/v29/i3/p518_

Beiser:1952:NET

[Bei52] Arthur Beiser. Nuclear emulsion technique. Reviews of Modern Physics, 24(4):273–311, October 1952. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.24.273;
 http://rmp.aps.org/abstract/RMP/v24/i4/p273_1.

Bell:1954:GTC

[Bel54] Dorothy G. Bell. Group theory and crystal lattices. Reviews of Modern Physics, 26(3):311–320, July 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.311;
 http://rmp.aps.org/abstract/RMP/v26/i3/p311_1.

Bennett:1959:FSC

[Ben59a] H. Stanley Bennett. Fine structure of cell nucleus, chromosomes, nucleoli, and membrane. Reviews of Modern Physics, 31(2):297–300, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.297;   http://
rmp.aps.org/abstract/RMP/v31/i2/p297_1.

Bennett:1959:SMC

[Ben59b] H. Stanley Bennett. Structure of muscle cells. Reviews of Modern Physics, 31(2):394–401, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.394;
 http://rmp.aps.org/abstract/RMP/v31/i2/p394_1.

Bergmann:1957:SCH

[Ber57] Peter G. Bergmann. Summary of the Chapel Hill Conference. Reviews of Modern Physics, 29(3):352–354, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.352;
 http://rmp.aps.org/abstract/RMP/v29/i3/p352_1.

Berliner:1959:MT

[Ber59] Robert W. Berliner. Membrane transport. Reviews of Modern Physics, 31(2):342–348, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.31.342;   http:/
/rmp.aps.org/abstract/RMP/v31/i2/p342_1.

Bethe:1950:RER

[Bet50] H. A. Bethe. The range-energy relation for slow alpha-particles and protons in air. Reviews of Modern Physics, 22(2):213–219, April 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.22.213;   http://rmp.aps.org/abstract/
RMP/v22/i2/p213_1.

Bethe:1955:MSH

[Bet55] Hans A. Bethe. Memorial Symposium Held in Honor of Enrico Fermi at the Washington Meeting of the American Physical Society, April 29, 1955. Reviews of Modern Physics, 27(3):249, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.27.249;   http://rmp.aps.org/abstract/
RMP/v27/i3/p249_1.

Biondi:1958:EEE

[BFGS58] M. A. Biondi, A. T. Forrester, M. P. Garfunkel, and C. B. Satterthwaite. Experimental evidence for an energy gap in superconductors. Reviews of Modern Physics, 30(4):1109–1136, October 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1109;   http://rmp.aps.org/abstract/RMP/
v30/i4/p1109_1.

Bernstein:1958:SN

[BG58] J. Bernstein and M. L. Goldberger. Structure of the nucleon. Reviews of Modern Physics, 30(2):465–470, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.465;
 http://rmp.aps.org/abstract/RMP/v30/i2/p465_1.

Berko:1956:ESP

[BH56] Stephan Berko and Frank L. Hereford. Experimental studies of positron interactions in solids and liquids. Reviews of Modern Physics, 28(3):299–307, July 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.299;
 http://rmp.aps.org/abstract/RMP/v28/i3/p299_1.

Bloch:1951:CFR

[BHB^{+}51] I. Bloch, M. H. Hull, Jr., A. A. Broyles, W. G. Bouricius, B. E. Freeman, and G. Breit. Coulomb functions for reactions of protons and alpha-particles with the lighter nuclei. Reviews of Modern Physics, 23(2):147–182, April 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.23.147;
 http://rmp.aps.org/abstract/RMP/v23/i2/p147_1.

Bickford:1953:LTT

[Bic53] L. R. Bickford, Jr. The low temperature transformation in ferrites. Reviews of Modern Physics, 25(1):75–79, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.75;   http://rmp.aps.org/abstract/RMP/v25/i1/p75_1.

Bade:1953:IS

[BJ53] W. L. Bade and Herbert Jehle. An introduction to spinors. Reviews of Modern Physics, 25(3):714–728, July 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.714;
 http://rmp.aps.org/abstract/RMP/v25/i3/p714_1.

Brix:1958:ISS

[BK58] Peter Brix and Hans Kopfermann. Isotope shift studies of nuclei. Reviews of Modern Physics, 30(2):517–520, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.517;   http://rmp.aps.org/abstract/RMP/v30/i2/p517_
1.

Bleaney:1953:APC

[Ble53a] B. Bleaney. Anomalous paramagnetism of copper acetate. Reviews of Modern Physics, 25(1):161–162, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.161;
 http://rmp.aps.org/abstract/RMP/v25/i1/p161_1.

Bleaney:1953:Db

[Ble53b] B. Bleaney. Discussion. Reviews of Modern Physics, 25 (1):232, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.232;   http://
rmp.aps.org/abstract/RMP/v25/i1/p232_1.

Bloom:1959:CR

[Blo59] William Bloom. Cellular responses. Reviews of Modern Physics, 31(1):21–29, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.31.21;   http://
rmp.aps.org/abstract/RMP/v31/i1/p21_1.

Bitter:1953:OIN

[BLR53] F. Bitter, R. F. Lacey, and B. Richter. Optically induced nuclear magnetization. A progress report. Reviews of Modern Physics, 25(1):174–176, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.174;
 http://rmp.aps.org/abstract/RMP/v25/i1/p174_1.

Bates:1953:HEM

[BM53] L. F. Bates and G. Marshall. Heat effects in the magnetization of silicon iron. Reviews of Modern Physics, 25 (1):17–33, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.17;   http://
rmp.aps.org/abstract/RMP/v25/i1/p17_1.

Bondi:1957:NMG

[Bon57] H. Bondi. Negative mass in General Relativity. Reviews of Modern Physics, 29(3):423–428, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.423;
 http://rmp.aps.org/abstract/RMP/v29/i3/p423_1.

Borgnis:1953:ARP

[Bor53] F. E. Borgnis. Acoustic radiation pressure of plane compressional waves. Reviews of Modern Physics, 25(3):653–664, July 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.653;   http://rmp.aps.org/abstract/
RMP/v25/i3/p653_1.

Bostick:1958:PHS

[Bos58] Winston H. Bostick. Possible hydromagnetic simulation of cosmical phenomena in the laboratory. Reviews of Modern Physics, 30(3):1090–1094, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.30.1090;   http://
rmp.aps.org/abstract/RMP/v30/i3/p1090_1.

Bozorth:1953:PP

[Boz53] R. M. Bozorth. The permalloy problem. Reviews of Modern Physics, 25(1):42–48, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.42;   http://
rmp.aps.org/abstract/RMP/v25/i1/p42_1.

Biedenharn:1953:TAC

[BR53] L. C. Biedenharn and M. E. Rose. Theory of angular correlation of nuclear radiations. Reviews of Modern Physics, 25(3):729–777, July 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.729;   http://rmp.aps.org/abstract/
RMP/v25/i3/p729_1.

Brysk:1958:TRO

[BR58] H. Brysk and M. E. Rose. Theoretical results on orbital capture. Reviews of Modern Physics, 30(4):1169–1177, October 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1169;   http://rmp.aps.org/abstract/RMP/
v30/i4/p1169_1. See [BR59].

Brysk:1959:ETR

[BR59] H. Brysk and M. E. Rose. Errata: Theoretical Results on Orbital Capture. Reviews of Modern Physics, 31(3):840, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.840;   http://rmp.aps.org/abstract/
RMP/v31/i3/p840_1. See [BR58].

Brattain:1951:COR

[Bra51] Walter H. Brattain. The copper oxide rectifier. Reviews of Modern Physics, 23(3):203–212, July 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.23.203;
 http://rmp.aps.org/abstract/RMP/v23/i3/p203_1.

Breit:1951:TST

[Bre51] G. Breit. Topics in scattering theory. Reviews of Modern Physics, 23(3):238–252, July 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.23.238;   http://
rmp.aps.org/abstract/RMP/v23/i3/p238_1.

Breene:1957:LS

[Bre57] R. G. Breene, Jr. Line shape. Reviews of Modern Physics, 29(1):94–143, January 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.29.94;   http://
rmp.aps.org/abstract/RMP/v29/i1/p94_1.

Breit:1958:TIS

[Bre58] G. Breit. Theory of isotope shift. Reviews of Modern Physics, 30(2):507–516, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.507;   http://
rmp.aps.org/abstract/RMP/v30/i2/p507_1.

Bridgman:1950:TPD

[Bri50] P. W. Bridgman. The thermodynamics of plastic deformation and generalized entropy. Reviews of Modern Physics, 22 (1):56–63, January 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.22.56;   http://
rmp.aps.org/abstract/RMP/v22/i1/p56_1.

Briggs:1954:ENA

[Bri54a] G. H. Briggs. The energies of natural alpha particles. Reviews of Modern Physics, 26(1):1–6, January 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.1;
http://rmp.aps.org/abstract/RMP/v26/i1/p1_1. See erratum

[Bri54b].

Briggs:1954:EEN

[Bri54b] G. H. Briggs. Erratum: The Energies of Natural Alpha Particles. Reviews of Modern Physics, 26(4):472, October 1954. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
26.472;   http://rmp.aps.org/abstract/RMP/v26/i4/p472_

1. See [Bri54a].

Brown:1953:MEF

[Bro53] William Fuller Brown, Jr. Magnetic energy formulas and their relation to magnetization theory. Reviews of Modern Physics, 25(1):131–135, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.131;   http://
rmp.aps.org/abstract/RMP/v25/i1/p131_1.

Brown:1959:FOM

[Bro59] G. E. Brown. Foundations of the optical model for nuclei and direct interaction. Reviews of Modern Physics, 31(4):893–919, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.893;   http://rmp.aps.org/abstract/
RMP/v31/i4/p893_1.

Barschall:1952:MFN

[BRTW52] H. H. Barschall, L. Rosen, R. F. Taschek, and J. H. Williams. Measurement of fast neutron flux. Reviews of Modern Physics, 24(1):1–14, January 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.24.1;   http://rmp.aps.
org/abstract/RMP/v24/i1/p1_1.

Brueckner:1958:RNS

[Bru58] K. A. Brueckner. Relation of nuclear size to saturation and nuclear compressibility. Reviews of Modern Physics, 30(2):561–565, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.561;   http://rmp.aps.org/abstract/
RMP/v30/i2/p561_1.

Blin-Stoyle:1956:TNM

[BS56] R. J. Blin-Stoyle. Theories of nuclear moments. Reviews of Modern Physics, 28(1):75–101, January 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
28.75;   http://rmp.aps.org/abstract/RMP/v28/i1/p75_1.

Bel:1958:GIM

[BS58a] N. Bel and E. Schatzman. On the gravitational instability of a medium in nonuniform rotation. Reviews of Modern Physics, 30(3):1015–1016, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.30.1015;   http://
rmp.aps.org/abstract/RMP/v30/i3/p1015_1.

Biermann:1958:MD

[BS58b] L. Biermann and A. Schlüter. Magnetohydrodynamic dissipation. Reviews of Modern Physics, 30(3):975–978, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.975;   http://rmp.aps.org/abstract/RMP/v30/i3/p975_
1.

Brockhouse:1958:ENM

[BS58c] B. N. Brockhouse and A. T. Stewart. Erratum: Normal Modes of Aluminum by Neutron Spectrometry. Reviews of Modern Physics, 30(4):1177, October 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.1177;   http:
//rmp.aps.org/abstract/RMP/v30/i4/p1177_1. See [BS58d].

Brockhouse:1958:NMA

[BS58d] B. N. Brockhouse and A. T. Stewart. Normal modes of aluminum by neutron spectrometry. Reviews of Modern Physics, 30(1):236–249, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.236;   http:/
/rmp.aps.org/abstract/RMP/v30/i1/p236_1.  See erratum

[BS58c].

Burgess:1958:CSP

[BS58e] A. Burgess and M. J. Seaton. Cross sections for photoionization from valence-electron states. Reviews of Modern Physics, 30(3):992–993, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.992;   http://
rmp.aps.org/abstract/RMP/v30/i3/p992_1.

Black:1958:ADE

[BT58a] P. J. Black and W. H. Taylor. Attempt to determine electron configurations in aluminum-rich alloys of transition metals. Reviews of Modern Physics, 30(1):55–58, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.55;   http://rmp.aps.org/abstract/RMP/v30/i1/p55_1.

Burgers:1958:P

[BT58b] J. M. Burgers and R. N. Thomas. Preface. Reviews of Modern Physics, 30(3):908–911, July 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.908;
 http://rmp.aps.org/abstract/RMP/v30/i3/p908_1.

Bullock:1959:INI

[Bul59] Theodore Holmes Bullock. Initiation of nerve impulses in receptor and central neurons. Reviews of Modern Physics, 31 (2):504–514, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.504;   http://
rmp.aps.org/abstract/RMP/v31/i2/p504_1.

Bass:1957:RCM

[BW57a] R. W. Bass and L. Witten. Remark on cosmological models. Reviews of Modern Physics, 29(3):452–453, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.452;
 http://rmp.aps.org/abstract/RMP/v29/i3/p452_1.

Brill:1957:ING

[BW57b] Dieter R. Brill and John A. Wheeler. Interaction of neutrinos and gravitational fields. Reviews of Modern Physics, 29(3):465–479, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.465;   http://rmp.aps.org/abstract/
RMP/v29/i3/p465_1. See errata [BW61].

Brill:1961:EIN

[BW61] Dieter R. Brill and John A. Wheeler. Errata: Interaction of Neutrinos and Gravitational Fields. Reviews of Modern Physics, 33 (4):623–624, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.33.623;   http://
rmp.aps.org/abstract/RMP/v33/i4/p623_2. See [BW57b].

Calvin:1959:ERT

[Cal59a] Melvin Calvin. Energy reception and transfer in photosynthesis. Reviews of Modern Physics, 31(1):147–156, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.147;   http://rmp.aps.org/abstract/
RMP/v31/i1/p147_1.

Calvin:1959:FRP

[Cal59b] Melvin Calvin. Free radicals in photosynthetic systems. Reviews of Modern Physics, 31(1):157–161, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.157;
 http://rmp.aps.org/abstract/RMP/v31/i1/p157_1.

Cameron:1958:SCF

[Cam58] A. G. W. Cameron. Semiempirical correlation of fission yield and kinetic energy distributions. Reviews of Modern Physics, 30(2):553–557, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.553;   http://
rmp.aps.org/abstract/RMP/v30/i2/p553_1.

Case:1957:TPR

[Cas57] K. M. Case. Transfer problems and the reciprocity principle. Reviews of Modern Physics, 29(4):651–663, October 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
29.651;   http://rmp.aps.org/abstract/RMP/v29/i4/p651_
1.

Cohen:1955:AVD

[CDLR55] E. Richard Cohen, Jesse W. DuMond, Thomas W. Layton, and John S. Rollett. Analysis of variance of the 1952 data on the atomic constants and a new adjustment, 1955. Reviews of Modern Physics, 27(4):363–380, October 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.27.363;
 http://rmp.aps.org/abstract/RMP/v27/i4/p363_1.

Carr:1953:D

[CE53] W. J. Carr, Jr. and H. Eckstein. Discussion. Reviews of Modern Physics, 25(1):107, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.107;
 http://rmp.aps.org/abstract/RMP/v25/i1/p107_1.

Crittenden:1953:TFF

[CH53] E. C. Crittenden, Jr. and R. W. Hoffman. Thin films of ferromagnetic materials. Reviews of Modern Physics, 25(1):310–315, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.310;   http://rmp.aps.org/abstract/
RMP/v25/i1/p310_1.

Cheng:1956:SDC

[Che56] L. S. Cheng. Sargent diagram and comparative half-lives for Kelectron capture transitions of odd nuclides (with Z ≤ 60). Reviews of Modern Physics, 28(1):45–47, January 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
28.45;   http://rmp.aps.org/abstract/RMP/v28/i1/p45_1.

Childs:1953:TEA

[Chi53] B. G. Childs. The thermal expansion of anisotropic metals. Reviews of Modern Physics, 25(3):665–670, July 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.665;
 http://rmp.aps.org/abstract/RMP/v25/i3/p665_1.

Clemence:1957:AT

[Cle57] G. M. Clemence. Astronomical time. Reviews of Modern Physics, 29(1):2–8, January 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.29.2;   http://rmp.aps.
org/abstract/RMP/v29/i1/p2_1.

Cochran:1958:SED

[Coc58] W. Cochran. Some electron distribution determinations by Xray diffraction. Reviews of Modern Physics, 30(1):47–50, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.47;   http://rmp.aps.org/abstract/RMP/
v30/i1/p47_1.

Cohen:1953:BCL

[Coh53] E. Richard Cohen. The basis for the criterion of least squares. Reviews of Modern Physics, 25(3):709–713, July 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.709;
 http://rmp.aps.org/abstract/RMP/v25/i3/p709_1.

Condon:1957:ICT

[Con57] E. U. Condon. International congress on theoretical physics. Reviews of Modern Physics, 29(2):159–160, April 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.159;
 http://rmp.aps.org/abstract/RMP/v29/i2/p159_1.

Connelly:1959:RPM

[Con59] C. M. Connelly. Recovery processes and metabolism of nerve. Reviews of Modern Physics, 31(2):475–484, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.475;
 http://rmp.aps.org/abstract/RMP/v31/i2/p475_1.

Cox:1950:SMG

[Cox50] R. T. Cox. The statistical method of Gibbs in irreversible change. Reviews of Modern Physics, 22(3):238–248, July 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
22.238;   http://rmp.aps.org/abstract/RMP/v22/i3/p238_
1.

Cox:1952:BMT

[Cox52] R. T. Cox. Brownian motion in the theory of irreversible processes. Reviews of Modern Physics, 24(4):312–320, October 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.24.312;   http://rmp.aps.org/abstract/
RMP/v24/i4/p312_1.

Cox:1958:CSB

[Cox58] E. G. Cox. Crystal structure of benzene. Reviews of Modern Physics, 30(1):159–162, January 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.159;
 http://rmp.aps.org/abstract/RMP/v30/i1/p159_1.

Cribier:1958:XRS

[Cri58] D. Cribier. X-ray scattering by thermal agitation and the Compton effect in fluorite. Reviews of Modern Physics, 30(1):228–229, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.228;   http://rmp.aps.org/abstract/
RMP/v30/i1/p228_1.

Cruickshank:1958:LVB

[Cru58] D. W. J. Cruickshank. On the lattice vibrations of benzene, naphthalene, and anthracene. Reviews of Modern Physics, 30 (1):163–167, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.163;   http://
rmp.aps.org/abstract/RMP/v30/i1/p163_1.

Curien:1958:CSX

[Cur58] H. Curien. Compton scattering of X-rays in crystals. Reviews of Modern Physics, 30(1):232–234, January 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.232;
 http://rmp.aps.org/abstract/RMP/v30/i1/p232_1.

Dalitz:1959:SPD

[Dal59] R. H. Dalitz. Strange particle decay processes and the Fermi interaction. Reviews of Modern Physics, 31(3):823–833, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.823;   http://rmp.aps.org/abstract/
RMP/v31/i3/p823_1.

Damon:1953:REF

[Dam53] R. W. Damon. Relaxation effects in the ferromagnetic resonance. Reviews of Modern Physics, 25(1):239–245, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.239;   http://rmp.aps.org/abstract/
RMP/v25/i1/p239_1.

Davies:1958:TOC

[Dav58] R. D. Davies. Twenty-one-cm studies of some interstellar clouds. Reviews of Modern Physics, 30(3):931–934, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.931;   http://rmp.aps.org/abstract/RMP/v30/i3/p931_
1.

Davisson:1953:NVR

[DB53] C. J. Davisson and J. W. Beams. A new variation of the rotationby-magnetization method of measuring gyromagnetic ratios. Reviews of Modern Physics, 25(1):246–252, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.246;   http://rmp.aps.org/abstract/RMP/v25/i1/p246_
1.

DuMond:1953:LSA

[DC53] Jesse W. M. DuMond and E. Richard Cohen. Least-squares adjustment of the atomic constants, 1952. Reviews of Modern Physics, 25(3):691–708, July 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.691;   http:/
/rmp.aps.org/abstract/RMP/v25/i3/p691_1.

Davisson:1952:GRA

[DE52] Charlotte Meaker Davisson and Robley D. Evans. Gammaray absorption coefficients. Reviews of Modern Physics, 24 (2):79–107, April 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.24.79;   http://
rmp.aps.org/abstract/RMP/v24/i2/p79_1.

Deser:1957:GRD

[Des57] S. Deser. General Relativity and the divergence problem in quantum field theory. Reviews of Modern Physics, 29(3):417–423, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.417;   http://rmp.aps.org/abstract/
RMP/v29/i3/p417_1.

DeWitt:1957:DTC

[DeW57a] Bryce S. DeWitt. Dynamical theory in curved spaces. I. A review of the classical and quantum action principles. Reviews of Modern Physics, 29(3):377–397, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.377;
 http://rmp.aps.org/abstract/RMP/v29/i3/p377_1.

DeWitt:1957:IN

[DeW57b] Bryce S. DeWitt. Introductory note. Reviews of Modern Physics, 29(3):351, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.351;   http://
rmp.aps.org/abstract/RMP/v29/i3/p351_1.

Duckworth:1954:MSA

[DHP54] Henry E. Duckworth, Benjamin G. Hogg, and Edwin M. Pennington. Mass spectroscopic atomic mass differences. Reviews of Modern Physics, 26(4):463–472, October 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.463;
 http://rmp.aps.org/abstract/RMP/v26/i4/p463_1.

Dicke:1957:GPE

[Dic57a] R. H. Dicke. Gravitation without a principle of equivalence. Reviews of Modern Physics, 29(3):363–376, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.363;
 http://rmp.aps.org/abstract/RMP/v29/i3/p363_1.

Dicke:1957:PEW

[Dic57b] R. H. Dicke. Principle of equivalence and the weak interactions. Reviews of Modern Physics, 29(3):355–362, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
29.355;   http://rmp.aps.org/abstract/RMP/v29/i3/p355_
1.

deKlerk:1953:NAD

[dKH53] D. de Klerk and R. P. Hudson. The new adiabatic demagnetization apparatus at the national bureau of standards. Reviews of Modern Physics, 25(1):292, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.292;
 http://rmp.aps.org/abstract/RMP/v25/i1/p292_1.

Dijkstra:1953:DPS

[DM53] L. J. Dijkstra and U. M. Martius. Domain pattern in silicon-iron under stress. Reviews of Modern Physics, 25(1):146–150, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.146;   http://rmp.aps.org/abstract/
RMP/v25/i1/p146_1.

Domenicali:1954:ITT

[Dom54] Charles A. Domenicali. Irreversible thermodynamics of thermoelectricity. Reviews of Modern Physics, 26(2):237–275, April 1954. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.26.237;   http://rmp.aps.org/abstract/
RMP/v26/i2/p237_1.

Donn:1958:VCM

[Don58] Bertram Donn. Validity of the cloud model of the interstellar medium. Reviews of Modern Physics, 30(3):940–942, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.940;   http://rmp.aps.org/abstract/
RMP/v30/i3/p940_1.

Doty:1959:CBI

[Dot59a] Paul Doty. Configurations of biologically important macromolecules in solution. Reviews of Modern Physics, 31(1): 107–117, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.107;   http://
rmp.aps.org/abstract/RMP/v31/i1/p107_1.

Doty:1959:PCM

[Dot59b] Paul Doty. Physicochemical characterization of macromolecules. Reviews of Modern Physics, 31(1):61–68, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
31.61;   http://rmp.aps.org/abstract/RMP/v31/i1/p61_1.

Daunt:1954:PLH

[DS54] J. G. Daunt and R. S. Smith. The problem of liquid helium-some recent aspects. Reviews of Modern Physics, 26(2):172–236, April 1954. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.26.172;   http://rmp.aps.org/abstract/
RMP/v26/i2/p172_1.

Duckworth:1957:MSA

[Duc57] Henry E. Duckworth. Mass spectroscopic atomic mass differences. II. Reviews of Modern Physics, 29(4):767–772, October 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.767;   http://rmp.aps.org/abstract/
RMP/v29/i4/p767_1.

deVaucouleurs:1958:LSS

[dV58] G. de Vaucouleurs. Large-scale structure and direction of rotation in galaxies. Reviews of Modern Physics, 30(3):926–930, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.926;   http://rmp.aps.org/abstract/
RMP/v30/i3/p926_1.

Endt:1957:ELL

[EB57] P. M. Endt and C. M. Braams. Energy levels of light nuclei ( Z =11 to Z =20 ). II. Reviews of Modern Physics, 29(4):683–756, October 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.683;   http://rmp.aps.org/abstract/
RMP/v29/i4/p683_1.

Eckart:1956:EMO

[EF56] Carl Eckart and Horace G. Ferris. Equations of motion of the ocean and atmosphere. Reviews of Modern Physics, 28 (1):48–52, January 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.28.48;   http://
rmp.aps.org/abstract/RMP/v28/i1/p48_1.

Eisenstein:1952:PPC

[Eis52] Julian Eisenstein. Physical properties of chromium alums at low temperatures. Reviews of Modern Physics, 24(2):74–78, April 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.24.74;   http://rmp.aps.org/abstract/RMP/
v24/i2/p74_1.

Eisenstein:1954:SE

[Eis54] Julian Eisenstein. Superconducting elements. Reviews of Modern Physics, 26(3):277–291, July 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.277;
 http://rmp.aps.org/abstract/RMP/v26/i3/p277_1.

Eisinger:1958:EDN

[EJ58] J. Eisinger and V. Jaccarino. Effects of the distribution of nuclear magnetization on hyperfine structure. Reviews of Modern Physics, 30(2):528–541, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.528;
 http://rmp.aps.org/abstract/RMP/v30/i2/p528_1.

Endt:1954:ELL

[EK54] P. M. Endt and J. C. Kluyver. Energy levels of light nuclei ( Z =11 to Z =20 ). Reviews of Modern Physics, 26 (1):95–166, January 1954. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.26.95;   http://
rmp.aps.org/abstract/RMP/v26/i1/p95_1.

Elliott:1953:CFT

[Ell53] R. J. Elliott. Crystal field theory in the rare earths. Reviews of Modern Physics, 25(1):167–169, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.167;
 http://rmp.aps.org/abstract/RMP/v25/i1/p167_1.

Ellas:1959:CIT

[Ell59] Peter Ellas. Coding and information theory. Reviews of Modern Physics, 31(1):221–226, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.221;
 http://rmp.aps.org/abstract/RMP/v31/i1/p221_1.

Elsasser:1950:EIG

[Els50] Walter M. Elsasser. The Earth’s interior and geomagnetism. Reviews of Modern Physics, 22(1):1–35, January 1950. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.22.1;
http://rmp.aps.org/abstract/RMP/v22/i1/p1_1.

Elsasser:1956:HDT

[Els56] Walter M. Elsasser. Hydromagnetic dynamo theory. Reviews of Modern Physics, 28(2):135–163, April 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.135;
 http://rmp.aps.org/abstract/RMP/v28/i2/p135_1.

Elton:1958:ISC

[Elt58] L. R. B. Elton. Interpretation of scattering cross sections in terms of nuclear size. Reviews of Modern Physics, 30(2):557–560, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.557;   http://rmp.aps.org/abstract/
RMP/v30/i2/p557_1.

Elliott:1958:TCS

[EM58] R. J. Elliott and W. Marshall. Theory of critical scattering. Reviews of Modern Physics, 30(1):75–89, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.75;   http://rmp.aps.org/abstract/RMP/v30/i1/p75_1.

Epstein:1956:TWP

[Eps56] Paul S. Epstein. Theory of wave propagation in a gyromagnetic medium. Reviews of Modern Physics, 28(1):3–17, January 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.28.3;   http://rmp.aps.org/
abstract/RMP/v28/i1/p3_1.

Ernst:1957:LTG

[Ern57] Frederick J. Ernst, Jr. Linear and toroidal geons. Reviews of Modern Physics, 29(3):496, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.496;
 http://rmp.aps.org/abstract/RMP/v29/i3/p496_1.

Everett:1957:RSF

[Eve57] Hugh Everett III. “relative state” formulation of quantum mechanics. Reviews of Modern Physics, 29(3):454–462, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.454;   http://rmp.aps.org/abstract/

RMP/v29/i3/p454_1. See assessment [Whe57].

Fano:1957:DSQ

[Fan57] U. Fano. Description of states in quantum mechanics by density matrix and operator techniques. Reviews of Modern Physics, 29(1):74–93, January 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.29.74;   http://
rmp.aps.org/abstract/RMP/v29/i1/p74_1.

Fletcher:1951:MRS

[FB51] C. H. Fletcher and Walker Bleakney. The Mach reflection of shock waves at nearly glancing incidence. Reviews of Modern Physics, 23(3):271–286, July 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.23.271;
 http://rmp.aps.org/abstract/RMP/v23/i3/p271_1.

Fowler:1956:MNT

[FB56] J. L. Fowler and John E. Brolley, Jr. Monoenergetic neutron techniques in the 10- to 30-MeV range. Reviews of Modern Physics, 28(2):103–134, April 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.103;
 http://rmp.aps.org/abstract/RMP/v28/i2/p103_1.

Feingold:1951:TVB

[Fei51] Arnold M. Feingold. Table of ft values in beta-decay. Reviews of Modern Physics, 23(1):10–20, January 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
23.10;   http://rmp.aps.org/abstract/RMP/v23/i1/p10_1.

Ferrell:1956:TPA

[Fer56] Richard A. Ferrell. Theory of positron annihilation in solids. Reviews of Modern Physics, 28(3):308–337, July 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.308;
 http://rmp.aps.org/abstract/RMP/v28/i3/p308_1.

Fernbach:1958:NRD

[Fer58] S. Fernbach. Nuclear radii as determined by scattering of neutrons. Reviews of Modern Physics, 30(2):414–418, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.414;   http://rmp.aps.org/abstract/RMP/v30/i2/p414_
1.

Ferry:1959:RMS

[Fer59] John D. Ferry. Rheology of macromolecular systems. Reviews of Modern Physics, 31(1):130–135, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.130;
 http://rmp.aps.org/abstract/RMP/v31/i1/p130_1.

Fessenkov:1958:SCD

[Fes58] V. G. Fessenkov. Star chains and dark filaments in galactic nebulae. Reviews of Modern Physics, 30(3):951–952, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.951;   http://rmp.aps.org/abstract/
RMP/v30/i3/p951_1.

Feynman:1957:SS

[Fey57] R. P. Feynman. Superfluidity and superconductivity. Reviews of Modern Physics, 29(2):205–212, April 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.205;
 http://rmp.aps.org/abstract/RMP/v29/i2/p205_1.

Fagg:1959:PMN

[FH59] Lawrence W. Fagg and Stanley S. Hanna. Polarization measurements on nuclear gamma rays. Reviews of Modern Physics, 31(3):711–758, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.711;   http://
rmp.aps.org/abstract/RMP/v31/i3/p711_1.

Fine:1953:MEA

[Fin53] M. E. Fine. Magnetomechanical effects in an antiferromagnet, CoO. Reviews of Modern Physics, 25(1):158, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.158;   http://rmp.aps.org/abstract/
RMP/v25/i1/p158_1.

Fernandez-Moran:1959:FSB

[FM59] H. Fernández-Morán. Fine structure of biological lamellar systems. Reviews of Modern Physics, 31(2):319–330, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
31.319;   http://rmp.aps.org/abstract/RMP/v31/i2/p319_
1.

Fock:1957:TLR

[Foc57] V. Fock. Three lectures on Relativity Theory. Reviews of Modern Physics, 29(3):325–333, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.325;
 http://rmp.aps.org/abstract/RMP/v29/i3/p325_1.

Foldy:1958:NEI

[Fol58] Leslie L. Foldy. Neutron-electron interaction. Reviews of Modern Physics, 30(2):471–481, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.471;
 http://rmp.aps.org/abstract/RMP/v30/i2/p471_1.

Froberg:1955:NTC

[Frö55] Carl-Erik Fröberg. Numerical treatment of Coulomb wave functions. Reviews of Modern Physics, 27(4):399–411, October 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.27.399;   http://rmp.aps.org/abstract/
RMP/v27/i4/p399_1.

Feenberg:1950:ICH

[FT50] Eugene Feenberg and George Trigg. The interpretation of comparative half-lives in the Fermi theory of beta-decay. Reviews of Modern Physics, 22(4):399–406, October 1950. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.22.399;
 http://rmp.aps.org/abstract/RMP/v22/i4/p399_1.

Feshbach:1958:ERW

[FV58] Herman Feshbach and Felix Villars. Elementary relativistic wave mechanics of spin 0 and spin 1/2 particles. Reviews of Modern Physics, 30(1):24–45, January 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.24;   http://rmp.aps.org/abstract/RMP/v30/i1/p24_1.

Fox:1953:D

[FW53] A. G. Fox and M. T. Weiss. Discussion. Reviews of Modern Physics, 25(1):262–263, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.262;
 http://rmp.aps.org/abstract/RMP/v25/i1/p262_1.

Gorter:1953:Da

[GA53] C. J. Gorter and A. Arrott. Discussion. Reviews of Modern Physics, 25(1):41, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.41;   http://
rmp.aps.org/abstract/RMP/v25/i1/p41_1.

Gabor:1956:TEI

[Gab56] D. Gabor. Theory of electron interference experiments. Reviews of Modern Physics, 28(3):260–276, July 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.260;
 http://rmp.aps.org/abstract/RMP/v28/i3/p260_1.

Galt:1953:MDW

[GAH53] J. K. Galt, J. Andrus, and H. G. Hopper. Motion of domain walls in ferrite crystals. Reviews of Modern Physics, 25 (1):93–97, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.93;   http://
rmp.aps.org/abstract/RMP/v25/i1/p93_1.

Ganguly:1959:ETO

[GC59] S. C. Ganguly and N. K. Chaudhury. Energy transport in organic phosphors. Reviews of Modern Physics, 31(4):990–1017, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.990;   http://rmp.aps.org/abstract/
RMP/v31/i4/p990_1.

Geisler:1953:CFP

[Gei53a] A. H. Geisler. Coercive force of precipitation alloys. Reviews of Modern Physics, 25(1):316–322, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.316;
 http://rmp.aps.org/abstract/RMP/v25/i1/p316_1.

Geisler:1953:D

[Gei53b] A. H. Geisler. Discussion. Reviews of Modern Physics, 25(1):74, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.74;   http://
rmp.aps.org/abstract/RMP/v25/i1/p74_1.

Gorter:1953:Db

[GG53] C. J. Gorter and J. E. Goldman. Discussion. Reviews of Modern Physics, 25(1):113, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.113;
 http://rmp.aps.org/abstract/RMP/v25/i1/p113_1.

Gilbert:1953:D

[GGHG53] T. L. Gilbert, C. J. Gorter, R. W. Hoffman, and C. Guillaud. Discussion. Reviews of Modern Physics, 25(1):315, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.315;   http://rmp.aps.org/abstract/
RMP/v25/i1/p315_1.

Geschwind:1954:DAM

[GGMT54] S. Geschwind, G. R. Gunther-Mohr, and C. H. Townes. Determination of atomic masses by microwave spectroscopy. Reviews of Modern Physics, 26(4):444–455, October 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.444;
 http://rmp.aps.org/abstract/RMP/v26/i4/p444_1.

Goldhaber:1952:NIS

[GH52] M. Goldhaber and R. D. Hill. Nuclear isomerism and shell structure. Reviews of Modern Physics, 24(3):179–239, July 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.24.179;   http://rmp.aps.org/abstract/
RMP/v24/i3/p179_1.

Goldman:1953:D

[GHS^{+}53] J. E. Goldman, J. M. Hastings, E. B. Stoner, L. R. Aronin, and G. W. Rathenau. Discussion. Reviews of Modern Physics, 25(1):53–55, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.53;   http://
rmp.aps.org/abstract/RMP/v25/i1/p53_1.

Glassgold:1958:NDD

[Gla58] A. E. Glassgold. Nuclear density distributions from proton scattering. Reviews of Modern Physics, 30(2):419–423, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.419;   http://rmp.aps.org/abstract/
RMP/v30/i2/p419_1.

Glimcher:1959:MBM

[Gli59] Melvin J. Glimcher. Molecular biology of mineralized tissues with particular reference to bone. Reviews of Modern Physics, 31(2):359–393, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.359;   http://
rmp.aps.org/abstract/RMP/v31/i2/p359_1.

Gell-Mann:1959:SWI

[GM59] Murray Gell-Mann. Status of weak interactions. Reviews of Modern Physics, 31(3):834–838, July 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.834;
 http://rmp.aps.org/abstract/RMP/v31/i3/p834_1.

Goldman:1953:AMA

[Gol53] J. E. Goldman. Atomic moments in alloys. Reviews of Modern Physics, 25(1):108–113, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.108;
 http://rmp.aps.org/abstract/RMP/v25/i1/p108_1.

Goldberg:1957:TGQ

[Gol57] Irwin Goldberg. Transformation groups in a Q-number configuration space. Reviews of Modern Physics, 29(3):450–451, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.450;   http://rmp.aps.org/abstract/
RMP/v29/i3/p450_1.

Goldsworthy:1958:PSI

[Gol58] F. A. Goldsworthy. On the propagation and structure of ionization fronts. Reviews of Modern Physics, 30(3):1062–1068, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1062;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1062_1.

Goldberger:1959:MCB

[Gol59] M. L. Goldberger. Mu capture, beta decay, and pi-meson decay. Reviews of Modern Physics, 31(3):797–801, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
31.797;   http://rmp.aps.org/abstract/RMP/v31/i3/p797_
1.

Good:1955:PDM

[Goo55] R. H. Good, Jr. Properties of the Dirac matrices. Reviews of Modern Physics, 27(2):187–211, April 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.27.187;
 http://rmp.aps.org/abstract/RMP/v27/i2/p187_1.

Gordon:1953:ESB

[Gor53a] D. I. Gordon. An experimental study of barkhausen noise in nickel-iron alloys. Reviews of Modern Physics, 25(1): 56–57, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.56;   http://
rmp.aps.org/abstract/RMP/v25/i1/p56_1.

Gorter:1953:AAR

[Gor53b] C. J. Gorter. Antiferromagnetism and antiferromagnetic resonance in a rhombic crystal at T =0. Reviews of Modern Physics, 25(1):277–278, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.277;
 http://rmp.aps.org/abstract/RMP/v25/i1/p277_1.

Gorter:1953:Dc

[Gor53c] C. J. Gorter. Discussion. Reviews of Modern Physics, 25 (1):165, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.165;   http://
rmp.aps.org/abstract/RMP/v25/i1/p165_2.

Gorter:1953:OAC

[Gor53d] C. J. Gorter. Observations on antiferromagnetic Cu Cl_{2} · 2 H_{2}O crystals. Reviews of Modern Physics, 25(1):332–337, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.332;   http://rmp.aps.org/abstract/
RMP/v25/i1/p332_1.

Grant:1959:PRT

[Gra59] F. A. Grant. Properties of rutile (titanium dioxide). Reviews of Modern Physics, 31(3):646–674, July 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.646;
 http://rmp.aps.org/abstract/RMP/v31/i3/p646_1.

Green:1958:NSW

[Gre58] Alex E. S. Green. Nuclear sizes and the Weiszäcker mass formula. Reviews of Modern Physics, 30(2):569–584, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.569;   http://rmp.aps.org/abstract/RMP/v30/i2/p569_
1.

Grayson-Smith:1953:D

[GS53] H. Grayson-Smith. Discussion. Reviews of Modern Physics, 25 (1):128–129, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.128;   http://
rmp.aps.org/abstract/RMP/v25/i1/p128_2.

Guillaud:1953:D

[GSB^{+}53] C. Guillaud, W. Sucksmith, L. F. Bates, C. Kittel, and E. C. Stoner. Discussion. Reviews of Modern Physics, 25 (1):308–309, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.308;   http://
rmp.aps.org/abstract/RMP/v25/i1/p308_1.

Guillaud:1953:EMM

[Gui53] Charles Guillaud. Elementary mechanisms of magnetization in mixed oxides of iron and cobalt. Reviews of Modern Physics, 25(1):64–73, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.64;   http://
rmp.aps.org/abstract/RMP/v25/i1/p64_1.

Gupta:1957:EOT

[Gup57] Suraj N. Gupta. Einstein’s and other theories of gravitation. Reviews of Modern Physics, 29(3):334–336, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.334;
 http://rmp.aps.org/abstract/RMP/v29/i3/p334_1.

Grenville-Wells:1958:EDD

[GW58] H. J. Grenville-Wells. Electron density distribution in crystals. Reviews of Modern Physics, 30(1):171–172, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.171;   http://rmp.aps.org/abstract/RMP/v30/i1/p171_
1.

Graham:1953:D

[GZ53] John W. Graham and Clarence Zener. Discussion. Reviews of Modern Physics, 25(1):295–296, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.295;
 http://rmp.aps.org/abstract/RMP/v25/i1/p295_1.

Haar:1950:CPS

[Haa50] D. Ter Haar. Cosmogonical problems and stellar energy. Reviews of Modern Physics, 22(2):119–152, April 1950. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.22.119;
 http://rmp.aps.org/abstract/RMP/v22/i2/p119_1.

Haar:1955:FSM

[Haa55] D. Ter Haar. Foundations of statistical mechanics. Reviews of Modern Physics, 27(3):289–338, July 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.27.289;
 http://rmp.aps.org/abstract/RMP/v27/i3/p289_1.

Haefner:1951:REL

[Hae51] Richard R. Haefner. Rotational energy levels of an alphaparticle model for the beryllium and carbon isotopes. Reviews of Modern Physics, 23(3):228–237, July 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.23.228;
 http://rmp.aps.org/abstract/RMP/v23/i3/p228_1.

Hagstrum:1951:IEI

[Hag51] Homer D. Hagstrum. Ionization by electron impact in CO, N_{2}, NO, and O_{2}. Reviews of Modern Physics, 23(3):185–203, July 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.23.185;   http://rmp.aps.org/abstract/
RMP/v23/i3/p185_1.

Hartree:1958:VAW

[Har58] D. R. Hartree. Variation of atomic wave functions with atomic number. Reviews of Modern Physics, 30(1):63–68, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.63;   http://rmp.aps.org/abstract/RMP/v30/
i1/p63_1.

Hartline:1959:RMI

[Har59] H. K. Hartline. Receptor mechanisms and the integration of sensory information in the eye. Reviews of Modern Physics, 31(2):515–523, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.515;   http://
rmp.aps.org/abstract/RMP/v31/i2/p515_1.

Hofstadter:1958:ESP

[HBY58] R. Hofstadter, F. Bumiller, and M. R. Yearian. Electromagnetic structure of the proton and neutron. Reviews of Modern Physics, 30(2):482–497, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.482;
 http://rmp.aps.org/abstract/RMP/v30/i2/p482_1.

Hastings:1953:NDS

[HC53] J. M. Hastings and L. M. Corliss. Neutron diffraction studies of zinc ferrite and nickel ferrite. Reviews of Modern Physics, 25(1):114–119, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.114;   http://
rmp.aps.org/abstract/RMP/v25/i1/p114_1.

Heisenberg:1957:QTF

[Hei57] W. Heisenberg. Quantum theory of fields and elementary particles. Reviews of Modern Physics, 29(3):269–278, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
29.269;   http://rmp.aps.org/abstract/RMP/v29/i3/p269_
1.

Heller:1959:ALS

[Hel59] Wilfried Heller. Anisotropic light scattering of streaming suspensions and solutions. Reviews of Modern Physics, 31(4): 1072–1077, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.31.1072;   http://
rmp.aps.org/abstract/RMP/v31/i4/p1072_1.

Henry:1953:SMS

[Hen53] Warren E. Henry. Some magnetization studies of Cr^{+++}, Fe^{+++}, Gd^{+++}, and Cu^{++} at low temperatures and in strong magnetic fields. Reviews of Modern Physics, 25(1):163–164, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.163;   http://rmp.aps.org/abstract/
RMP/v25/i1/p163_1.

Henley:1958:NRM

[Hen58] Ernest M. Henley. Nuclear radii from mesonic atoms. Reviews of Modern Physics, 30(2):438–441, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.438;
 http://rmp.aps.org/abstract/RMP/v30/i2/p438_1.

Herman:1958:TIE

[Her58] Frank Herman. Theoretical investigation of the electronic energy band structure of solids. Reviews of Modern Physics, 30 (1):102–121, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.102;   http://
rmp.aps.org/abstract/RMP/v30/i1/p102_1.

Hess:1958:SHE

[Hes58] Wilmot N. Hess. Summary of high-energy nucleon-nucleon crosssection data. Reviews of Modern Physics, 30(2):368–401, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.368;   http://rmp.aps.org/abstract/
RMP/v30/i2/p368_1.

Hill:1951:HPC

[Hil51] E. L. Hill. Hamilton’s Principle and the conservation theorems of mathematical physics. Reviews of Modern Physics, 23(3):253–260, July 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.23.253;   http://rmp.aps.org/abstract/
RMP/v23/i3/p253_1.

Hulthen:1951:AED

[HL51] Lamek Hulthén and K. V. Laurikainen. Approximate eigensolutions of (d^{2}ϕ/dx^{2}) + [a + b(e^{−x}/x)]ϕ = 0. Reviews of Modern Physics, 23(1):1–9, January 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.23.1;
http://rmp.aps.org/abstract/RMP/v23/i1/p1_1.

Hershberger:1953:PRP

[HL53] W. D. Hershberger and H. N. Leifer. Paramagnetic resonance in phosphors at microwave frequencies. Reviews of Modern Physics, 25(1):291, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.291;   http:/
/rmp.aps.org/abstract/RMP/v25/i1/p291_1.

Hornyak:1950:ELL

[HLMF50] W. F. Hornyak, T. Lauritsen, P. Morrison, and W. A. Fowler. Energy levels of light nuclei. III. Reviews of Modern Physics, 22(4):291–372, October 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.22.291;   http://
rmp.aps.org/abstract/RMP/v22/i4/p291_1.

Hodge:1959:FPM

[Hod59a] Alan J. Hodge. Fibrous proteins of muscle. Reviews of Modern Physics, 31(2):409–425, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.409;
 http://rmp.aps.org/abstract/RMP/v31/i2/p409_1.

Hodge:1959:FSL

[Hod59b] Alan J. Hodge. Fine structure of lamellar systems as illustrated by chloroplasts. Reviews of Modern Physics, 31(2):331–341, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.331;   http://rmp.aps.org/abstract/
RMP/v31/i2/p331_1.

Hofstadter:1956:ESN

[Hof56] Robert Hofstadter. Electron scattering and nuclear structure. Reviews of Modern Physics, 28(3):214–254, July 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.214;
 http://rmp.aps.org/abstract/RMP/v28/i3/p214_1.

Hofstadter:1958:ICN

[Hof58] R. Hofstadter. International Congress on Nuclear Sizes and Density Distributions Held at Stanford University, December 17– 19, 1957. Reviews of Modern Physics, 30(2):412–413, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.412;   http://rmp.aps.org/abstract/
RMP/v30/i2/p412_1.

Hogan:1953:FFE

[Hog53] C. L. Hogan. The ferromagnetic Faraday effect at microwave frequencies and its applications. Reviews of Modern Physics, 25 (1):253–262, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.253;   http://
rmp.aps.org/abstract/RMP/v25/i1/p253_1.

Hogness:1959:IES

[Hog59] David S. Hogness. Induced enzyme synthesis. Reviews of Modern Physics, 31(1):256–268, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.256;
 http://rmp.aps.org/abstract/RMP/v31/i1/p256_1.

Hutchison:1953:PRA

[HP53] Clyde A. Hutchison, Jr. and Ricardo C. Pastor. Paramagnetic resonance absorption in solutions of K in liquid NH_{3}. Reviews of Modern Physics, 25(1):285–290, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.285;
 http://rmp.aps.org/abstract/RMP/v25/i1/p285_1.

Hollander:1953:TI

[HPS53] J. M. Hollander, I. Perlman, and G. T. Seaborg. Table of isotopes. Reviews of Modern Physics, 25(2):469–651, April 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.469;   http://rmp.aps.org/abstract/RMP/v25/i2/p469_
1.

Infeld:1951:FM

[IH51] L. Infeld and T. E. Hull. The factorization method. Reviews of Modern Physics, 23(1):21–68, January 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
23.21;   http://rmp.aps.org/abstract/RMP/v23/i1/p21_1.

Infeld:1957:EMG

[Inf57] Leopold Infeld. Equations of motion in General Relativity theory and the action principle. Reviews of Modern Physics, 29(3):398–411, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.398;   http://
rmp.aps.org/abstract/RMP/v29/i3/p398_1.

Inglis:1953:ELS

[Ing53] D. R. Inglis. The energy levels and the structure of light nuclei. Reviews of Modern Physics, 25(2):390–450, April 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.390;
 http://rmp.aps.org/abstract/RMP/v25/i2/p390_1. See erra-

tum [Ing55a].

Inglis:1955:EEL

[Ing55a] D. R. Inglis. Erratum: The Energy Levels and the Structure of Light Nuclei. Reviews of Modern Physics, 27(1):76, January 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.27.76;   http://rmp.aps.org/abstract/RMP/v27/

i1/p76_1. See [Ing53].

Inglis:1955:TEM

[Ing55b] D. R. Inglis. Theories of the Earth’s magnetism. Reviews of Modern Physics, 27(2):212–248, April 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.27.212;
 http://rmp.aps.org/abstract/RMP/v27/i2/p212_1.

Inglis:1957:SEA

[Ing57] David R. Inglis. Shifting of the Earth’s axis of rotation. Reviews of Modern Physics, 29(1):9–19, January 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.9;
http://rmp.aps.org/abstract/RMP/v29/i1/p9_1.

Inoue:1959:MCM

[Ino59] Shinya Inoué. Motility of cilia and the mechanism of mitosis. Reviews of Modern Physics, 31(2):402–408, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.402;
 http://rmp.aps.org/abstract/RMP/v31/i2/p402_1.

Jacobsen:1958:LDT

[Jac58] E. H. Jacobsen. Lattice dynamics and thermal diffuse scattering of X-rays. Reviews of Modern Physics, 30(1):234–235, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.234;   http://rmp.aps.org/abstract/
RMP/v30/i1/p234_1.

Jackson:1950:ILE

[JB50] J. David Jackson and John M. Blatt. The interpretation of low energy proton-proton scattering. Reviews of Modern Physics, 22(1):77–118, January 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.22.77;   http://
rmp.aps.org/abstract/RMP/v22/i1/p77_1.

Jensen:1957:NSM

[Jen57] J. H. D. Jensen. Nuclear shell models. Reviews of Modern Physics, 29(2):182–185, April 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.29.182;   http:/
/rmp.aps.org/abstract/RMP/v29/i2/p182_1.

Joseph:1958:PPB

[JR58] J. Joseph and F. Rohrlich. Pair production and Bremsstrahlung in the field of free and bound electrons. Reviews of Modern Physics, 30(2):354–368, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.354;   http:/
/rmp.aps.org/abstract/RMP/v30/i2/p354_1.  See erratum

[JR59].

Joseph:1959:EPP

[JR59] J. Joseph and F. Rohrlich. Errata: Pair Production and Bremsstrahlung in the Field of Free and Bound Electrons. Reviews of Modern Physics, 31(3):840, July 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
31.840;   http://rmp.aps.org/abstract/RMP/v31/i3/p840_

2. See [JR58].

Kittel:1953:RPF

[KA53] C. Kittel and Elihu Abrahams. Relaxation process in ferromagnetism. Reviews of Modern Physics, 25(1):233–238, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.233;   http://rmp.aps.org/abstract/
RMP/v25/i1/p233_1.

Kahn:1958:CTH

[Kah58a] F. D. Kahn. Collision of two highly ionized clouds of gas. Reviews of Modern Physics, 30(3):1069–1072, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1069;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1069_1.

Kahn:1958:SIF

[Kah58b] F. D. Kahn. On the stability of ionization fronts. Reviews of Modern Physics, 30(3):1058–1061, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1058;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1058_1.

Kaplan:1958:FIG

[Kap58a] S. A. Kaplan. On the formation of interstellar gas clouds. Reviews of Modern Physics, 30(3):943, July 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.943;
 http://rmp.aps.org/abstract/RMP/v30/i3/p943_1.

Kaplan:1958:SWM

[Kap58b] S. A. Kaplan. Shock waves in magnetogasodynamic turbulence. Reviews of Modern Physics, 30(3):1089, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1089;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1089_1.

Kasha:1959:RBE

[Kas59] Michael Kasha. Relation between exciton bands and conduction bands in molecular lamellar systems. Reviews of Modern Physics, 31(1):162–169, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.162;
 http://rmp.aps.org/abstract/RMP/v31/i1/p162_1.

Katz:1959:MST

[Kat59a] Bernard Katz. Mechanisms of synaptic transmission. Reviews of Modern Physics, 31(2):524–531, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.524;
 http://rmp.aps.org/abstract/RMP/v31/i2/p524_1.

Katz:1959:NNI

[Kat59b] Bernard Katz. Nature of the nerve impulse. Reviews of Modern Physics, 31(2):466–474, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.466;
 http://rmp.aps.org/abstract/RMP/v31/i2/p466_1.

Kauzmann:1959:CSB

[Kau59] Walter Kauzmann. Chemical specificity in biological systems. Reviews of Modern Physics, 31(2):549–556, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.549;
 http://rmp.aps.org/abstract/RMP/v31/i2/p549_1.

Kaya:1953:UAP

[Kay53] Seiji Kaya. Uniaxial anisotropy of a permalloy crystal. Reviews of Modern Physics, 25(1):49–53, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.49;   http://rmp.aps.org/abstract/RMP/v25/i1/p49_1.

Kahl:1958:CRI

[KB58] G. D. Kahl and F. D. Bennett. Coherence requirements for interferometry. Reviews of Modern Physics, 30(4):1193– 1196, October 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.30.1193;   http://
rmp.aps.org/abstract/RMP/v30/i4/p1193_1.

Kendrew:1959:TDS

[Ken59] John C. Kendrew. Three-dimensional structure of globular proteins. Reviews of Modern Physics, 31(1):94–99, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.31.94;   http://rmp.aps.org/abstract/RMP/v31/
i1/p94_1.

Kerr:1958:SWC

[Ker58] F. J. Kerr. Sydney work on 21-cm observations of the HI galactic disk. Reviews of Modern Physics, 30(3):924–925, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.924;   http://rmp.aps.org/abstract/
RMP/v30/i3/p924_1.

Kittel:1953:D

[KG53] C. Kittel and C. P. Guillaud. Discussion. Reviews of Modern Physics, 25(1):119–121, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.119;
 http://rmp.aps.org/abstract/RMP/v25/i1/p119_1.

Kofoed-Hansen:1958:MND

[KH58] O. Kofoed-Hansen. Mirror nuclei determinations of nuclear size. Reviews of Modern Physics, 30(2):449–456, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.449;   http://rmp.aps.org/abstract/RMP/v30/i2/p449_
1.

Kihara:1952:MTE

[Kih52] Taro Kihara. The mathematical theory of electrical discharges in gases. Reviews of Modern Physics, 24(1):45–61, January 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.24.45;   http://rmp.aps.org/abstract/RMP/v24/
i1/p45_1.

Kihara:1953:MTE

[Kih53a] Taro Kihara. The mathematical theory of electrical discharges in gases. B. Velocity-distribution of positive ions in a static field. Reviews of Modern Physics, 25(4):844–852, October 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.844;   http://rmp.aps.org/abstract/RMP/v25/i4/p844_
1.

Kihara:1953:VCM

[Kih53b] Taro Kihara. Virial coefficients and models of molecules in gases. Reviews of Modern Physics, 25(4):831–843, October 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.831;   http://rmp.aps.org/abstract/
RMP/v25/i4/p831_1.

Kihara:1955:VCM

[Kih55] Taro Kihara. Virial coefficients and models of molecules in gases. B. Reviews of Modern Physics, 27(4):412–423, October 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.27.412;   http://rmp.aps.org/abstract/
RMP/v27/i4/p412_1.

King:1954:TTB

[Kin54] R. W. King. Table of total beta-disintegration energies. Reviews of Modern Physics, 26(4):327–401, October 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.327;
 http://rmp.aps.org/abstract/RMP/v26/i4/p327_1.

Kip:1953:MRA

[Kip53] Arthur F. Kip. Microwave resonance absorption in gadolinium metal. Reviews of Modern Physics, 25(1):229–231, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.229;   http://rmp.aps.org/abstract/
RMP/v25/i1/p229_1.

Kittel:1953:SEI

[Kit53] C. Kittel. Symposium on exchange-introduction. Reviews of Modern Physics, 25(1):191, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.191;
 http://rmp.aps.org/abstract/RMP/v25/i1/p191_1.

Klinkenberg:1952:TNS

[Kli52] P. F. A. Klinkenberg. Tables of nuclear shell structure. Reviews of Modern Physics, 24(2):63–73, April 1952. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
24.63;   http://rmp.aps.org/abstract/RMP/v24/i2/p63_1.

Knopoff:1958:ASA

[KM58] Leon Knopoff and Gordon J. F. MacDonald. Attenuation of small amplitude stress waves in solids. Reviews of Modern Physics, 30 (4):1178–1192, October 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.30.1178;   http://
rmp.aps.org/abstract/RMP/v30/i4/p1178_1.

Koch:1959:BCS

[KM59] H. W. Koch and J. W. Motz. Bremsstrahlung cross-section formulas and related data. Reviews of Modern Physics, 31 (4):920–955, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.920;   http://
rmp.aps.org/abstract/RMP/v31/i4/p920_1.

Kohn:1954:ECB

[Koh54a] Walter Kohn. Erratum: On the Convergence of Born Expansions. Reviews of Modern Physics, 26(4):472, October 1954. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
26.472;   http://rmp.aps.org/abstract/RMP/v26/i4/p472_

2. See [Koh54b].

Kohn:1954:CBE

[Koh54b] Walter Kohn. On the convergence of Born expansions. Reviews of Modern Physics, 26(3):292–310, July 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.292;
 http://rmp.aps.org/abstract/RMP/v26/i3/p292_1. See erra-

tum [Koh54a].

Konopinski:1955:FTB

[Kon55] E. J. Konopinski. Fermi’s theory of beta-decay. Reviews of Modern Physics, 27(3):254–257, July 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.27.254;
 http://rmp.aps.org/abstract/RMP/v27/i3/p254_1.

Kornberg:1959:BNA

[Kor59] Arthur Kornberg. Biosynthesis of nucleic acids. Reviews of Modern Physics, 31(1):200–209, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.200;
 http://rmp.aps.org/abstract/RMP/v31/i1/p200_1.

Katz:1952:RER

[KP52] L. Katz and A. S. Penfold. Range-energy relations for electrons and the determination of beta-ray end-point energies by absorption. Reviews of Modern Physics, 24(1):28–44, January 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
24.28;   http://rmp.aps.org/abstract/RMP/v24/i1/p28_1.

Karim:1952:SCV

[KR52] S. M. Karim and L. Rosenhead. The second coefficient of viscosity of liquids and gases. Reviews of Modern Physics, 24 (2):108–116, April 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.24.108;   http://
rmp.aps.org/abstract/RMP/v24/i2/p108_1.

Kriessman:1953:HTM

[Kri53] C. J. Kriessman. The high temperature magnetic susceptibility of V, Nb, Ta, W, and Mo. Reviews of Modern Physics, 25 (1):122–126, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.122;   http://
rmp.aps.org/abstract/RMP/v25/i1/p122_1.

Kubo:1953:SWT

[Kub53] Ryogo Kubo. The spin-wave theory as a variational method and its application to antiferromagnetism. Reviews of Modern Physics, 25(1):344–351, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.344;
 http://rmp.aps.org/abstract/RMP/v25/i1/p344_1.

Kursunoglu:1957:CGT

[Kur57] Behram Kursunoglu. Correspondence in the generalized theory of gravitation. Reviews of Modern Physics, 29(3):412–416, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.412;   http://rmp.aps.org/abstract/
RMP/v29/i3/p412_1.

Koehler:1953:PSN

[KW53] W. C. Koehler and E. O. Wollan. Paramagnetic scattering of neutrons by trivalent ions of Nd and Er. Reviews of Modern Physics, 25(1):128, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.128;
 http://rmp.aps.org/abstract/RMP/v25/i1/p128_1.

Klauder:1957:QNA

[KW57] John Klauder and John A. Wheeler. On the question of a neutrino analog to electric charge. Reviews of Modern Physics, 29(3):516–517, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.516;   http://
rmp.aps.org/abstract/RMP/v29/i3/p516_1.

Landsberg:1956:FT

[Lan56] P. T. Landsberg. Foundations of thermodynamics. Reviews of Modern Physics, 28(4):363–392, October 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.363;
 http://rmp.aps.org/abstract/RMP/v28/i4/p363_1.

Lanczos:1957:EGR

[Lan57a] Cornelius Lanczos. Electricity and General Relativity. Reviews of Modern Physics, 29(3):337–350, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.337;
 http://rmp.aps.org/abstract/RMP/v29/i3/p337_1.

Lane:1957:NR

[Lan57b] A. M. Lane. Nuclear reactions. Reviews of Modern Physics, 29(2):191–199, April 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.191;   http://
rmp.aps.org/abstract/RMP/v29/i2/p191_1.

Laval:1958:XRS

[Lav58] J. Laval. X-ray scattering by thermal agitation of atoms in a crystal. Reviews of Modern Physics, 30(1):222–227, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.222;   http://rmp.aps.org/abstract/
RMP/v30/i1/p222_1.

Lax:1951:MSW

[Lax51] Melvin Lax. Multiple scattering of waves. Reviews of Modern Physics, 23(4):287–310, October 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.23.287;
 http://rmp.aps.org/abstract/RMP/v23/i4/p287_1.

Lax:1958:EIE

[Lax58] Benjamin Lax. Experimental investigations of the electronic band structure of solids. Reviews of Modern Physics, 30(1): 122–154, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.122;   http://
rmp.aps.org/abstract/RMP/v30/i1/p122_1.

LeCorre:1958:EPE

[Le 58] Y. Le Corre. Electrical polarization of elastic waves in crystals and optical dispersion. Reviews of Modern Physics, 30 (1):229–231, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.229;   http://
rmp.aps.org/abstract/RMP/v30/i1/p229_1.

Lehninger:1959:RET

[Leh59] Albert L. Lehninger. Respiratory-energy transformation. Reviews of Modern Physics, 31(1):136–146, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.136;
 http://rmp.aps.org/abstract/RMP/v31/i1/p136_1.

Lennox:1959:GFS

[Len59] E. S. Lennox. Genetic fine-structure analysis. Reviews of Modern Physics, 31(1):242–248, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.242;
 http://rmp.aps.org/abstract/RMP/v31/i1/p242_1.

Levinthal:1959:CAP

[Lev59a] Cyrus Levinthal. Coding aspects of protein synthesis. Reviews of Modern Physics, 31(1):249–255, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.249;
 http://rmp.aps.org/abstract/RMP/v31/i1/p249_1.

Levinthal:1959:GRN

[Lev59b] Cyrus Levinthal. Genetics and the role of nucleic acids. Reviews of Modern Physics, 31(1):227–232, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.227;
 http://rmp.aps.org/abstract/RMP/v31/i1/p227_1.

Lewis:1952:MMP

[Lew52] H. W. Lewis. Multiple meson production in nucleon-nucleon collisions. Reviews of Modern Physics, 24(4):241–248, October 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.24.241;   http://rmp.aps.org/abstract/
RMP/v24/i4/p241_1.

Lidofsky:1957:TTB

[Lid57] L. J. Lidofsky. Table of total beta-disintegration energies. II. Reviews of Modern Physics, 29(4):773–830, October 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
29.773;   http://rmp.aps.org/abstract/RMP/v29/i4/p773_
1.

Lax:1953:D

[LK53a] M. Lax and F. Keffer. Discussion. Reviews of Modern Physics, 25(1):337, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.337;   http://
rmp.aps.org/abstract/RMP/v25/i1/p337_1.

Lin:1953:HCP

[LK53b] S. T. Lin and A. R. Kaufmann. Helmholtz coils for production of powerful and uniform fields and gradients. Reviews of Modern Physics, 25(1):182–190, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.182;
 http://rmp.aps.org/abstract/RMP/v25/i1/p182_1.

Lonsdale:1958:TVA

[Lon58] Kathleen Lonsdale. Thermal vibrations of atoms and molecules in crystals. Reviews of Modern Physics, 30(1):168–170, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.168;   http://rmp.aps.org/abstract/
RMP/v30/i1/p168_1.

Low:1957:PM

[Low57] F. E. Low. Pi mesodynamics. Reviews of Modern Physics, 29(2):216–222, April 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.216;   http://
rmp.aps.org/abstract/RMP/v29/i2/p216_1.

Lowde:1958:CMS

[Low58] R. D. Lowde. Critical magnetic scattering of neutrons by iron. Reviews of Modern Physics, 30(1):69–74, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.69;   http://rmp.aps.org/abstract/RMP/v30/i1/p69_1.

Levy:1958:NDS

[LP58] H. A. Levy and S. W. Peterson. Neutron diffraction studies of hydrogen containing crystals. Reviews of Modern Physics, 30(1):101, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.101;   http://
rmp.aps.org/abstract/RMP/v30/i1/p101_1.

Leiss:1958:NMD

[LS58] J. E. Leiss and R. A. Schrack. Nuclear matter distributions from coherent neutral pion production. Reviews of Modern Physics, 30(2):456–461, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.456;   http://
rmp.aps.org/abstract/RMP/v30/i2/p456_1.

Lin:1959:IRM

[LS59] Chun C. Lin and Jerome D. Swalen. Internal rotation and microwave spectroscopy. Reviews of Modern Physics, 31(4): 841–892, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.841;   http://
rmp.aps.org/abstract/RMP/v31/i4/p841_1.

Legvold:1953:SME

[LSBE53] S. Legvold, F. H. Spedding, F. Barson, and J. F. Elliott. Some magnetic and electrical properties of gadolinium, dysprosium, and erbium metals. Reviews of Modern Physics, 25(1):129–130, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.129;   http://rmp.aps.org/abstract/
RMP/v25/i1/p129_1.

Levy:1953:UAM

[LT53] Sheldon Levy and Rohn Truell. Ultrasonic attenuation in magnetic single crystals. Reviews of Modern Physics, 25(1):140–145, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.140;   http://rmp.aps.org/abstract/
RMP/v25/i1/p140_1.

Lane:1958:RMT

[LT58] A. M. Lane and R. G. Thomas. R-matrix theory of nuclear reactions. Reviews of Modern Physics, 30(2):257–353, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.257;   http://rmp.aps.org/abstract/
RMP/v30/i2/p257_1.

Lindquist:1957:DLU

[LW57] Richard W. Lindquist and John A. Wheeler. Dynamics of a lattice universe by the schwarzschild-cell method. Reviews of Modern Physics, 29(3):432–443, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.432;
 http://rmp.aps.org/abstract/RMP/v29/i3/p432_1. See erra-

tum [LW59].

Lindquist:1959:EDL

[LW59] Richard W. Lindquist and John A. Wheeler. Erratum: Dynamics of a Lattice Universe by the Schwarzschild-Cell Method. Reviews of Modern Physics, 31(3):839, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.31.839;   http:/
/rmp.aps.org/abstract/RMP/v31/i3/p839_4. See [LW57].

Ma:1953:IVL

[Ma53] S. T. Ma. Interpretation of the virtual level of the deuteron. Reviews of Modern Physics, 25(4):853–860, October 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.853;   http://rmp.aps.org/abstract/RMP/v25/i4/p853_
1.

Mack:1950:TNM

[Mac50] J. E. Mack. A table of nuclear moments, January 1950. Reviews of Modern Physics, 22(1):64–76, January 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
22.64;   http://rmp.aps.org/abstract/RMP/v22/i1/p64_1.

Marshak:1951:MRH

[Mar51] R. E. Marshak. Meson reactions in hydrogen and deuterium. Reviews of Modern Physics, 23(2):137–146, April 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.23.137;
 http://rmp.aps.org/abstract/RMP/v23/i2/p137_1.

Marton:1956:CQI

[Mar56a] L. Marton. Conference on quantum interactions of the free electron. Reviews of Modern Physics, 28(3):171, July 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
28.171;   http://rmp.aps.org/abstract/RMP/v28/i3/p171_

Marton:1956:ELE

[Mar56b] L. Marton. Experiments on low-energy electron scattering and energy losses. Reviews of Modern Physics, 28(3):172–183, July 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.28.172;   http://rmp.aps.org/abstract/
RMP/v28/i3/p172_1.

Markham:1959:INM

[Mar59] Jordan J. Markham. Interaction of normal modes with electron traps. Reviews of Modern Physics, 31(4):956–989, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.956;   http://rmp.aps.org/abstract/
RMP/v31/i4/p956_1.

Mason:1953:RRN

[Mas53] W. P. Mason. Rotational relaxation in nickel at high frequencies. Reviews of Modern Physics, 25(1):136–139, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.136;   http://rmp.aps.org/abstract/
RMP/v25/i1/p136_1.

Massey:1956:TSS

[Mas56] H. S. W. Massey. Theory of the scattering of slow electrons. Reviews of Modern Physics, 28(3):199–213, July 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.199;
 http://rmp.aps.org/abstract/RMP/v28/i3/p199_1.

Major:1954:SDC

[MB54] J. K. Major and L. C. Biedenharn. Sargent diagram and comparative half-lives for electron capture transitions. Reviews of Modern Physics, 26(3):321–326, July 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.321;
 http://rmp.aps.org/abstract/RMP/v26/i3/p321_1.

Macdonald:1956:LSI

[MB56] J. Ross Macdonald and Malcolm K. Brachman. Linear-system integral transform relations. Reviews of Modern Physics, 28 (4):393–422, October 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.28.393;   http://
rmp.aps.org/abstract/RMP/v28/i4/p393_1.

Markham:1951:ASF

[MBL51] Jordan J. Markham, Robert T. Beyer, and R. B. Lindsay. Absorption of sound in fluids. Reviews of Modern Physics, 23 (4):353–411, October 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.23.353;   http://
rmp.aps.org/abstract/RMP/v23/i4/p353_1.

McVittie:1958:SES

[McV58] G. C. McVittie. Some exact solutions of the equations of magnetohydrodynamics when both self-attraction and magnetic fields are present. Reviews of Modern Physics, 30(3):1080–1083, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1080;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1080_1.

Mrozowski:1953:D

[ME53] S. Mrozowski and W. P. Eatherly. Discussion. Reviews of Modern Physics, 25(1):180–181, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.180;
 http://rmp.aps.org/abstract/RMP/v25/i1/p180_1.

Megaw:1958:SEP

[Meg58] Helen D. Megaw. Some empirical problems concerning atomic bonds. Reviews of Modern Physics, 30(1):96–99, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.96;   http://rmp.aps.org/abstract/RMP/v30/
i1/p96_1.

Meiklejohn:1953:ESC

[Mei53] W. H. Meiklejohn. Experimental study of the coercive force of fine particles. Reviews of Modern Physics, 25(1):302–306, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.25.302;   http://rmp.aps.org/abstract/
RMP/v25/i1/p302_1.

Meister:1959:ABP

[Mei59] Alton Meister. Approaches to the biosynthesis of proteins. Reviews of Modern Physics, 31(1):210–220, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.210;
 http://rmp.aps.org/abstract/RMP/v31/i1/p210_1.

Melvin:1956:SFS

[Mel56] M. A. Melvin. Simplification in finding symmetry-adapted eigenfunctions. Reviews of Modern Physics, 28(1):18–44, January 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.28.18;   http://rmp.aps.org/abstract/RMP/v28/
i1/p18_1.

Mestel:1958:MFC

[Mes58] L. Mestel. Magnetic field of a contracting protostar. Reviews of Modern Physics, 30(3):1020–1022, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1020;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1020_1.

Miyake:1958:TCX

[MH58] Shizuo Miyake and Sadao Hoshino. Temperature characteristic of X-ray intensity reflection from crystals having zincblendeand wurtzite-type structures. Reviews of Modern Physics, 30 (1):172–174, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.172;   http://
rmp.aps.org/abstract/RMP/v30/i1/p172_1.

Michel:1957:WIB

[Mic57] Louis Michel. Weak interactions between “old” particles and beta decay. Reviews of Modern Physics, 29(2):223–230, April 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.223;   http://rmp.aps.org/abstract/
RMP/v29/i2/p223_1.

Milburn:1955:STM

[Mil55] Richard H. Milburn. Statistical theory of multiple meson production. Reviews of Modern Physics, 27(1):1–14, January 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.27.1;   http://rmp.aps.org/
abstract/RMP/v27/i1/p1_1.

Minkowski:1958:CLS

[Min58] R. Minkowski. Cygnus loop and some related nebulosities. Reviews of Modern Physics, 30(3):1048–1052, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1048;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1048_1.

Misner:1957:FQG

[Mis57] Charles W. Misner. Feynman quantization of General Relativity. Reviews of Modern Physics, 29(3):497–509, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
29.497;   http://rmp.aps.org/abstract/RMP/v29/i3/p497_
1.

Mitchell:1950:SSA

[Mit50] Allan C. G. Mitchell. Spectroscopy of some artificially radioactive nuclei. Reviews of Modern Physics, 22(1):36–55, January 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.22.36;   http://rmp.aps.org/abstract/RMP/v22/
i1/p36_1.

Margenau:1959:SSL

[ML59] Henry Margenau and Marvin Lewis. Structure of spectral lines from plasmas. Reviews of Modern Physics, 31(3):569–615, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.569;   http://rmp.aps.org/abstract/
RMP/v31/i3/p569_1.

Maxwell:1953:AR

[MM53] L. R. Maxwell and T. R. McGuire. Antiferromagnetic resonance. Reviews of Modern Physics, 25(1):279–284, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.279;   http://rmp.aps.org/abstract/
RMP/v25/i1/p279_1.

Maradudin:1958:RVD

[MMMW58] A. A. Maradudin, P. Mazur, E. W. Montroll, and G. H. Weiss. Remarks on the vibrations of diatomic lattices. Reviews of Modern Physics, 30(1):175–196, January 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.175;
 http://rmp.aps.org/abstract/RMP/v30/i1/p175_1.

Mayer:1951:NSS

[MMN51] M. G. Mayer, S. A. Moszkowski, and L. W. Nordheim. Nuclear shell structure and beta-decay. I. Odd A nuclei. Reviews of Modern Physics, 23(4):315–321, October 1951. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.23.315;
 http://rmp.aps.org/abstract/RMP/v23/i4/p315_1.

Morrison:1957:OCR

[Mor57] Philip Morrison. On the origins of cosmic rays. Reviews of Modern Physics, 29(2):235–243, April 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.235;
 http://rmp.aps.org/abstract/RMP/v29/i2/p235_1.

Morales:1959:MMC

[Mor59] Manuel F. Morales. Mechanisms of muscle contraction. Reviews of Modern Physics, 31(2):426–432, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.426;
 http://rmp.aps.org/abstract/RMP/v31/i2/p426_1.

Mottelson:1957:CMN

[Mot57] Ben R. Mottelson. Collective motion in the nucleus. Reviews of Modern Physics, 29(2):186–190, April 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.186;
 http://rmp.aps.org/abstract/RMP/v29/i2/p186_1.

Matthews:1951:RMT

[MS51] P. T. Matthews and Abdus Salam. The renormalization of meson theories. Reviews of Modern Physics, 23(4):311–314, October 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.23.311;   http://rmp.aps.org/abstract/
RMP/v23/i4/p311_1.

Munch:1958:IMO

[Mün58a] G. Münch. Internal motions in the orion nebula. Reviews of Modern Physics, 30(3):1035–1041, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1035;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1035_1.

Munch:1958:KFC

[Mün58b] G. Münch. Kinematics of the filaments in the crab nebula. Reviews of Modern Physics, 30(3):1042–1046, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1042;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1042_1.

Miles:1957:DSF

[MWV57] P. A. Miles, W. B. Westphal, and A. Von Hippel. Dielectric spectroscopy of ferromagnetic semiconductors. Reviews of Modern Physics, 29(3):279–307, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.279;
 http://rmp.aps.org/abstract/RMP/v29/i3/p279_1.

Myers:1952:DIa

[Mye52a] W. R. Myers. The diamagnetism of ions. Reviews of Modern Physics, 24(1):15–27, January 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.24.15;   http://
rmp.aps.org/abstract/RMP/v24/i1/p15_1.

Myers:1952:DIb

[Mye52b] W. R. Myers. The diamagnetism of ions. Reviews of Modern Physics, 24(3):178, July 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.24.178;   http://
rmp.aps.org/abstract/RMP/v24/i3/p178_1.

Newman:1957:OST

[NB57] Ezra Newman and Peter G. Bergmann. Observables in singular theories by systematic approximation. Reviews of Modern Physics, 29(3):443–449, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.29.443;   http:/
/rmp.aps.org/abstract/RMP/v29/i3/p443_1.

Neel:1953:SNR

[Née53a] Louis Néel. Some new results on antiferromagnetism and ferromagnetism. Reviews of Modern Physics, 25(1):58–63, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.25.58;   http://rmp.aps.org/abstract/RMP/v25/
i1/p58_1.

Neel:1953:TMF

[Née53b] Louis Néel. Thermoremanent magnetization of fine powders. Reviews of Modern Physics, 25(1):293–295, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.293;   http://rmp.aps.org/abstract/RMP/v25/i1/p293_
1.

Neurath:1959:PSE

[Neu59] Hans Neurath. Protein structure and enzyme action. Reviews of Modern Physics, 31(1):185–190, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.185;
 http://rmp.aps.org/abstract/RMP/v31/i1/p185_1.

Nesbitt:1953:PMS

[NH53] E. A. Nesbitt and R. D. Heidenreich. The physical and magnetic structure of the mishima alloys. Reviews of Modern Physics, 25(1):322–323, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.322;
 http://rmp.aps.org/abstract/RMP/v25/i1/p322_1.

Nielsen:1951:VRE

[Nie51] Harald H. Nielsen. The vibration-rotation energies of molecules. Reviews of Modern Physics, 23(2):90–136, April 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
23.90;   http://rmp.aps.org/abstract/RMP/v23/i2/p90_1.

Newell:1953:TIMa

[NM53a] Gordon F. Newell and Elliott W. Montroll. On the theory of the Ising model of ferromagnetism. Reviews of Modern Physics, 25(1):159, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.159;   http://
rmp.aps.org/abstract/RMP/v25/i1/p159_1.

Newell:1953:TIMb

[NM53b] Gordon F. Newell and Elliott W. Montroll. On the theory of the Ising model of ferromagnetism. Reviews of Modern Physics, 25(2):353–389, April 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.353;   http://
rmp.aps.org/abstract/RMP/v25/i2/p353_1.

Nagamiya:1958:TMS

[NM58] T. Nagamiya and K. Motizuki. Theory of the magnetic scattering of neutrons by CoO. Reviews of Modern Physics, 30 (1):89–93, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.89;   http://
rmp.aps.org/abstract/RMP/v30/i1/p89_1.

Nordheim:1951:NSS

[Nor51] L. W. Nordheim. Nuclear shell structure and beta-decay. II. Even A nuclei. Reviews of Modern Physics, 23(4):322–327, October 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.23.322;   http://rmp.aps.org/abstract/
RMP/v23/i4/p322_1.

Nussbaum:1956:ESP

[Nus56] Rudi H. Nussbaum. Experimental single-particle level spacings in the region of the 1f_{7/2} shell. Reviews of Modern Physics, 28(4):423–431, October 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.28.423;   http://
rmp.aps.org/abstract/RMP/v28/i4/p423_1.

Oncley:1959:CCP

[Onc59] J. L. Oncley. Chemical characterization of proteins, carbohydrates, and lipids. Reviews of Modern Physics, 31(1): 30–49, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.31.30;   http://
rmp.aps.org/abstract/RMP/v31/i1/p30_1.

Opechowski:1953:MOE

[Ope53] W. Opechowski. Magneto-optical effects and paramagnetic resonance. Reviews of Modern Physics, 25(1):264–268, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.264;   http://rmp.aps.org/abstract/
RMP/v25/i1/p264_1.

Oppenheimer:1956:E

[Opp56] J. Robert Oppenheimer. Einstein. Reviews of Modern Physics, 28(1):1–2, January 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.28.1;   http://rmp.aps.
org/abstract/RMP/v28/i1/p1_1. In memoriam tribute.

Orgel:1959:HB

[Org59] Leslie E. Orgel. The hydrogen bond. Reviews of Modern Physics, 31(1):100–102, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.100;
 http://rmp.aps.org/abstract/RMP/v31/i1/p100_1.

Odgers:1958:PMS

[OS58] G. J. Odgers and R. W. Stewart. Possible mechanisms of single and multiple star formation. Reviews of Modern Physics, 30(3):1017–1019, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.30.1017;   http://
rmp.aps.org/abstract/RMP/v30/i3/p1017_1.

Opechowski:1953:D

[OWL53] W. Opechowski, G. H. Wannier, and M. Lax. Discussion. Reviews of Modern Physics, 25(1):159–160, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.159;
 http://rmp.aps.org/abstract/RMP/v25/i1/p159_2.

Page:1959:EPP

[Pag59] Lorne A. Page. Electron and positron polarization. Reviews of Modern Physics, 31(3):759–781, July 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.759;
 http://rmp.aps.org/abstract/RMP/v31/i3/p759_1.

Parker:1958:GDI

[Par58] E. N. Parker. Gross dynamics of the interstellar medium. Reviews of Modern Physics, 30(3):955–965, July 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.955;
 http://rmp.aps.org/abstract/RMP/v30/i3/p955_1.

Parratt:1959:EBS

[Par59] L. G. Parratt. Electronic band structure of solids by X-ray spectroscopy. Reviews of Modern Physics, 31(3):616–645, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.616;   http://rmp.aps.org/abstract/
RMP/v31/i3/p616_1.

Pepinsky:1958:SNX

[Pep58] Ray Pepinsky. Some new X-ray and neutron studies of hydrogen bonding. Reviews of Modern Physics, 30(1):100, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.100;   http://rmp.aps.org/abstract/
RMP/v30/i1/p100_1.

Petschek:1958:AD

[Pet58] H. E. Petschek. Aerodynamic dissipation. Reviews of Modern Physics, 30(3):966–974, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.966;   http:/
/rmp.aps.org/abstract/RMP/v30/i3/p966_1.

Pines:1956:CEL

[Pin56] David Pines. Collective energy losses in solids. Reviews of Modern Physics, 28(3):184–198, July 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.184;
 http://rmp.aps.org/abstract/RMP/v28/i3/p184_1.

Pollard:1959:RIE

[Pol59] Ernest Pollard. Radiation inactivation of enzymes, nucleic acids, and phage particles. Reviews of Modern Physics, 31 (2):273–281, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.273;   http://
rmp.aps.org/abstract/RMP/v31/i2/p273_1.

Post:1956:CFR

[Pos56] Richard F. Post. Controlled fusion research — an application of the physics of high temperature plasmas. Reviews of Modern Physics, 28(3):338–362, July 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.338;
 http://rmp.aps.org/abstract/RMP/v28/i3/p338_1.

Pottasch:1958:BRD

[Pot58] Stuart R. Pottasch. Bright rims in diffuse nebulae. Reviews of Modern Physics, 30(3):1053–1058, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1053;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1053_1.

Pugh:1953:HEF

[PR53] Emerson M. Pugh and Norman Rostoker. Hall effect in ferromagnetic materials. Reviews of Modern Physics, 25(1):151–157, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.151;   http://rmp.aps.org/abstract/
RMP/v25/i1/p151_1.

Pringsheim:1952:JF

[Pri52] Peter Pringsheim. James franck. Reviews of Modern Physics, 24(3):117–119, July 1952. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.24.117;   http://
rmp.aps.org/abstract/RMP/v24/i3/p117_1.

Primakoff:1959:TMC

[Pri59] H. Primakoff. Theory of muon capture. Reviews of Modern Physics, 31(3):802–822, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.31.802;   http:/
/rmp.aps.org/abstract/RMP/v31/i3/p802_1.

Polder:1953:RPF

[PS53] D. Polder and J. Smit. Resonance phenomena in ferrites. Reviews of Modern Physics, 25(1):89–90, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.89;   http://rmp.aps.org/abstract/RMP/v25/i1/p89_1.

Pickelner:1958:NGH

[PS58] S. B. Pickelner and I. S. Shklovsky. On the nature of the galactic halo. Reviews of Modern Physics, 30(3):935–939, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.935;   http://rmp.aps.org/abstract/
RMP/v30/i3/p935_1.

Puck:1959:QSM

[Puc59] T. T. Puck. Quantitative studies on mammalian cells in vitro. Reviews of Modern Physics, 31(2):433–448, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.433;
 http://rmp.aps.org/abstract/RMP/v31/i2/p433_1.

Power:1957:TG

[PW57] Edwin A. Power and John A. Wheeler. Thermal geons. Reviews of Modern Physics, 29(3):480–495, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.480;
 http://rmp.aps.org/abstract/RMP/v29/i3/p480_1.

Rado:1953:Da

[Rad53a] G. T. Rado. Discussion. Reviews of Modern Physics, 25 (1):90, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.90;   http://
rmp.aps.org/abstract/RMP/v25/i1/p90_1.

Rado:1953:Db

[Rad53b] G. T. Rado. Discussion. Reviews of Modern Physics, 25 (1):92, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.92;   http://
rmp.aps.org/abstract/RMP/v25/i1/p92_1.

Rado:1953:MSF

[Rad53c] George T. Rado. Magnetic spectra of ferrites. Reviews of Modern Physics, 25(1):81–89, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.81;   http://rmp.aps.org/abstract/RMP/v25/i1/p81_1.

Rasmussen:1958:NRA

[Ras58] John O. Rasmussen. Nuclear radii from alpha-particle scattering and alpha-radioactivity. Reviews of Modern Physics, 30(2):424–429, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.424;   http://
rmp.aps.org/abstract/RMP/v30/i2/p424_1.

Rathenau:1953:SMH

[Rat53] G. W. Rathenau. Saturation and magnetization of hexagonal iron oxide compounds. Reviews of Modern Physics, 25(1):297–301, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.297;   http://rmp.aps.org/abstract/
RMP/v25/i1/p297_1.

Ravenhall:1958:ESN

[Rav58] D. G. Ravenhall. Electron scattering and nuclear charge distributions. Reviews of Modern Physics, 30(2):430–438, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.430;   http://rmp.aps.org/abstract/
RMP/v30/i2/p430_1.

Rathenau:1953:D

[RBFF53] G. Rathenau, L. F. Bates, C. A. Fowler, Jr., and E. M. Fryer. Discussion. Reviews of Modern Physics, 25(1):33, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.25.33;   http://rmp.aps.org/abstract/RMP/v25/
i1/p33_1.

Robinson:1955:KCR

[RF55] Berol L. Robinson and Richard W. Fink. L / K-capture ratios, mean L-fluorescence yields, and transition energies in orbital electron-capture. Reviews of Modern Physics, 27(4): 424–430, October 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.27.424;   http://
rmp.aps.org/abstract/RMP/v27/i4/p424_1.

Rathenau:1953:Db

[RHWR53] G. W. Rathenau, C. L. Hogan, M. T. Weiss, and G. W. Rathenau. Discussion. Reviews of Modern Physics, 25(1):301, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.301;   http://rmp.aps.org/abstract/
RMP/v25/i1/p301_1.

Richardson:1955:AVS

[Ric55] E. G. Richardson. Absorption and velocity of sound in vapors. Reviews of Modern Physics, 27(1):15–25, January 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
27.15;   http://rmp.aps.org/abstract/RMP/v27/i1/p15_1.

Rice:1959:P

[Ric59a] Stuart A. Rice. Polyelectrolytes. Reviews of Modern Physics, 31(1):69–83, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.31.69;   http://
rmp.aps.org/abstract/RMP/v31/i1/p69_1.

Rich:1959:MCS

[Ric59b] Alexander Rich. Molecular configuration of synthetic and biological polymers. Reviews of Modern Physics, 31(1):50–60, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.50;   http://rmp.aps.org/abstract/RMP/
v31/i1/p50_1.

Rich:1959:MSN

[Ric59c] Alexander Rich. Molecular structure of the nucleic acids. Reviews of Modern Physics, 31(1):191–199, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.191;
 http://rmp.aps.org/abstract/RMP/v31/i1/p191_1.

Rado:1953:Dc

[RK53] G. T. Rado and C. Kittel. Discussion. Reviews of Modern Physics, 25(1):238, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.238;   http:/
/rmp.aps.org/abstract/RMP/v25/i1/p238_1.

Robertson:1958:SNA

[Rob58] J. Monteath Robertson. Structure of naphthalene and anthracene. Reviews of Modern Physics, 30(1):155–158, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.155;   http://rmp.aps.org/abstract/
RMP/v30/i1/p155_1.

Roberts:1959:GPB

[Rob59] Richard Brooke Roberts. General patterns of biochemical synthesis. Reviews of Modern Physics, 31(1):170–176, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.170;   http://rmp.aps.org/abstract/
RMP/v31/i1/p170_1.

Rojkovsky:1958:SSG

[Roj58] D. A. Rojkovsky. Statistical study of globules projected on stellar clouds. Reviews of Modern Physics, 30(3):949–950, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.949;   http://rmp.aps.org/abstract/
RMP/v30/i3/p949_1.

Roothaan:1951:NDM

[Roo51] C. C. J. Roothaan. New developments in molecular orbital theory. Reviews of Modern Physics, 23(2):69–89, April 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
23.69;   http://rmp.aps.org/abstract/RMP/v23/i2/p69_1.

Rosenblith:1959:SPO

[Ros59a] Walter A. Rosenblith. Sensory performance of organisms. Reviews of Modern Physics, 31(2):485–491, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.485;
 http://rmp.aps.org/abstract/RMP/v31/i2/p485_1.

Rosenblith:1959:SQA

[Ros59b] Walter A. Rosenblith. Some quantifiable aspects of the electrical activity of the nervous system (with emphasis upon responses to sensory stimuli). Reviews of Modern Physics, 31(2):532–545, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.532;   http://rmp.aps.org/abstract/
RMP/v31/i2/p532_1.

Rabi:1954:URC

[RRS54] I. I. Rabi, N. F. Ramsey, and J. Schwinger. Use of rotating coordinates in magnetic resonance problems. Reviews of Modern Physics, 26(2):167–171, April 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.167;
 http://rmp.aps.org/abstract/RMP/v26/i2/p167_1.

Rosen:1957:GFA

[RS57] Nathan Rosen and Hadassah Shamir. Gravitational field of an axially symmetric system in first approximation. Reviews of Modern Physics, 29(3):429–431, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.429;
 http://rmp.aps.org/abstract/RMP/v29/i3/p429_1.

Roberts:1953:MNE

[RSB53] L. D. Roberts, C. C. Sartain, and B. Borie. Measurement of the nuclear and electronic contributions to the specific heat of neodymium ethyl sulfate near 1 ^{∘}K. Reviews of Modern Physics, 25(1):170–173, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.170;
 http://rmp.aps.org/abstract/RMP/v25/i1/p170_1.

Rosenblum:1958:RDA

[RTG58] B. Rosenblum, C. H. Townes, and S. Geschwind. Recent determinations of atomic mass ratios by microwave spectroscopy. Reviews of Modern Physics, 30(2):409–411, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.409;
 http://rmp.aps.org/abstract/RMP/v30/i2/p409_1.

Safronov:1958:TPC

[Saf58] V. S. Safronov. On the turbulence in the protoplanetary cloud. Reviews of Modern Physics, 30(3):1023–1024, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1023;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1023_1.

Salpeter:1957:SES

[Sal57] E. E. Salpeter. Stellar energy sources. Reviews of Modern Physics, 29(2):244–254, April 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.244;   http://
rmp.aps.org/abstract/RMP/v29/i2/p244_1.

Savedoff:1958:GDE

[Sav58a] M. P. Savedoff. Gas-dynamical effects of star formation. Reviews of Modern Physics, 30(3):1073–1074, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1073;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1073_1.

Savedoff:1958:GDG

[Sav58b] M. P. Savedoff. Gas dynamics of galaxy collisions. Reviews of Modern Physics, 30(3):1074–1076, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.30.1074;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1074_1.

Stewart:1958:VSV

[SB58] A. T. Stewart and B. N. Brockhouse. Vibration spectra of vanadium and a Mn–Co alloy by neutron spectrometry. Reviews of Modern Physics, 30(1):250–255, January 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.250;
 http://rmp.aps.org/abstract/RMP/v30/i1/p250_1.

Scheidegger:1953:GM

[Sch53] A. E. Scheidegger. Gravitational motion. Reviews of Modern Physics, 25(2):451–468, April 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.451;   http:/
/rmp.aps.org/abstract/RMP/v25/i2/p451_1.

Schatzman:1958:CIMb

[Sch58a] E. Schatzman. On the condensation of interstellar material. Reviews of Modern Physics, 30(3):1012–1014, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1012;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1012_1.

Schatzman:1958:CIMa

[Sch58b] E. Schatzman. On the cooling of interstellar matter. Reviews of Modern Physics, 30(3):1009–1011, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1009;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1009_1.

Schiff:1958:ESN

[Sch58c] L. I. Schiff. Electromagnetic structure of the neutron. Reviews of Modern Physics, 30(2):462–464, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.462;
 http://rmp.aps.org/abstract/RMP/v30/i2/p462_1.

Schmitt:1959:IPE

[Sch59a] Francis O. Schmitt. Interaction properties of elongate protein macromolecules with particular reference to collagen (tropocollagen). Reviews of Modern Physics, 31(2):349–358, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.349;   http://rmp.aps.org/abstract/
RMP/v31/i2/p349_1.

Schmitt:1959:MBP

[Sch59b] Francis O. Schmitt. Molecular biology and the physical basis of life processes. Reviews of Modern Physics, 31(1):5–10, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.31.5;   http://rmp.aps.org/
abstract/RMP/v31/i1/p5_1.

Schmitt:1959:MON

[Sch59c] Francis O. Schmitt. Molecular organization of the nerve fiber. Reviews of Modern Physics, 31(2):455–465, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.455;
 http://rmp.aps.org/abstract/RMP/v31/i2/p455_1.

Schmitt:1959:BTC

[Sch59d] Otto H. Schmitt. Biological transducers and coding. Reviews of Modern Physics, 31(2):492–503, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.492;
 http://rmp.aps.org/abstract/RMP/v31/i2/p492_1.

Sucksmith:1953:SMT

[SCOT53] W. Sucksmith, C. A. Clark, D. J. Oliver, and J. E. Thompson. Spontaneous magnetization — techniques and measurements. Reviews of Modern Physics, 25(1):34–41, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.34;   http://rmp.aps.org/abstract/RMP/v25/i1/p34_1.

Seaton:1958:LDV

[Sea58a] M. J. Seaton. Local density variations in planetary nebulae. Reviews of Modern Physics, 30(3):1034, July 1958. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1034;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1034_1.

Seaton:1958:TIC

[Sea58b] M. J. Seaton. Thermal inelastic collision processes. Reviews of Modern Physics, 30(3):979–991, July 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.979;
 http://rmp.aps.org/abstract/RMP/v30/i3/p979_1.

Sedov:1958:EGM

[Sed58] L. I. Sedov. Examples of gas motion and certain hypotheses on the mechanism of stellar outbursts. Reviews of Modern Physics, 30(3):1077–1079, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.30.1077;   http://
rmp.aps.org/abstract/RMP/v30/i3/p1077_1.

Segre:1955:FNP

[Seg55] Emilio Segrè. Fermi and neutron physics. Reviews of Modern Physics, 27(3):257–263, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.27.257;   http:/
/rmp.aps.org/abstract/RMP/v27/i3/p257_1.

Segre:1958:NRA

[Seg58] E. Segrè. Nuclear radii from antiproton measurements. Reviews of Modern Physics, 30(2):550–552, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.550;
 http://rmp.aps.org/abstract/RMP/v30/i2/p550_1.

Seitz:1951:SPS

[Sei51] Frederick Seitz. Speculations on the properties of the silver halide crystals. Reviews of Modern Physics, 23(4):328–352, October 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.23.328;   http://rmp.aps.org/abstract/
RMP/v23/i4/p328_1.

Seitz:1954:CCA

[Sei54] Frederick Seitz. Color centers in alkali halide crystals. II. Reviews of Modern Physics, 26(1):7–94, January 1954. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.26.7;
http://rmp.aps.org/abstract/RMP/v26/i1/p7_1.

Seitz:1955:FSA

[Sei55] Frederick Seitz. Fermi statistics and its applications. Reviews of Modern Physics, 27(3):249–253, July 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.27.249;
 http://rmp.aps.org/abstract/RMP/v27/i3/p249_2.

Serkowski:1958:MGM

[Ser58] K. Serkowski. Microstructure of the galactic magnetic field. Reviews of Modern Physics, 30(3):952–954, July 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.952;
 http://rmp.aps.org/abstract/RMP/v30/i3/p952_1.

Seth:1958:NRL

[Set58] Kamal K. Seth. Nuclear radii by low-energy neutron scattering. Reviews of Modern Physics, 30(2):442–449, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.442;   http://rmp.aps.org/abstract/RMP/v30/i2/p442_
1.

Segre:1949:NI

[SH49] Emilio Segrè and A. C. Helmholz. Nuclear isomerism. Reviews of Modern Physics, 21(2):271–304, April 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.271;
 http://rmp.aps.org/abstract/RMP/v21/i2/p271_1. See erra-

tum [SH50].

Segre:1950:ENI

[SH50] Emilio Segré and A. C. Helmholz. Erratum: Nuclear Isomerism. Reviews of Modern Physics, 22(3):237, July 1950. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.

22.237;   http://rmp.aps.org/abstract/RMP/v22/i3/p237_

1. See [SH49].

Shapiro:1956:TPE

[Sha56] Anatole M. Shapiro. Table of properties of the “elementary” particles. Reviews of Modern Physics, 28(2):164–170, April 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.28.164;   http://rmp.aps.org/abstract/
RMP/v28/i2/p164_1.

Shacklett:1958:NSE

[Sha58] Robert L. Shacklett. Nuclear size estimates from X-ray fine structure measurements. Reviews of Modern Physics, 30 (2):521–527, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.521;   http://
rmp.aps.org/abstract/RMP/v30/i2/p521_1.

Shklovsky:1958:NEC

[Shk58] I. S. Shklovsky. On the nature of the emission of the crab nebula. Reviews of Modern Physics, 30(3):1047, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1047;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1047_1.

Strominger:1958:TI

[SHS58] D. Strominger, J. M. Hollander, and G. T. Seaborg. Table of isotopes. Reviews of Modern Physics, 30(2):585–904, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.585;   http://rmp.aps.org/abstract/
RMP/v30/i2/p585_1.

Simpson:1956:EIE

[Sim56] J. Arol Simpson. Electron interference experiments. Reviews of Modern Physics, 28(3):254–260, July 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.254;
 http://rmp.aps.org/abstract/RMP/v28/i3/p254_1.

Sjostrand:1959:FSC

[Sjö59] Fritiof S. Sjöstrand. Fine structure of cytoplasm: The organization of membranous layers. Reviews of Modern Physics, 31(2):301–318, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.301;   http://
rmp.aps.org/abstract/RMP/v31/i2/p301_1.

Slater:1953:FBT

[Sla53] J. C. Slater. Ferromagnetism and the band theory. Reviews of Modern Physics, 25(1):199–210, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.199;
 http://rmp.aps.org/abstract/RMP/v25/i1/p199_1.

Slater:1958:ICC

[Sla58a]  J. C. Slater.  International conference on current prob-

lems in crystal physics. Reviews of Modern Physics, 30 (1):46, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.46;   http://
rmp.aps.org/abstract/RMP/v30/i1/p46_1.

Slater:1958:IWC

[Sla58b] John C. Slater. Interaction of waves in crystals. Reviews of Modern Physics, 30(1):197–222, January 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.197;
 http://rmp.aps.org/abstract/RMP/v30/i1/p197_1.

Slepian:1960:CRH

[Sle60] Joseph Slepian. Comment on “Review of High-Temperature Rotating-Plasma Experiments”. Reviews of Modern Physics, 32 (4):1032–1033, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.1032;   http://
rmp.aps.org/abstract/RMP/v32/i4/p1032_1. See [Wil59a].

Stout:1953:MAI

[SM53] J. W. Stout and L. M. Matarrese. Magnetic anisotropy of the irongroup fluorides. Reviews of Modern Physics, 25(1):338–343, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.338;   http://rmp.aps.org/abstract/
RMP/v25/i1/p338_1.

Smart:1953:MFT

[Sma53] J. Samuel Smart. The molecular field treatment of antiferromagnetism. Reviews of Modern Physics, 25(1):327–331, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.327;   http://rmp.aps.org/abstract/
RMP/v25/i1/p327_1.

Shankland:1955:NAI

[SMLK55] R. S. Shankland, S. W. McCuskey, F. C. Leone, and G. Kuerti. New analysis of the interferometer observations of dayton C. Miller. Reviews of Modern Physics, 27(2):167–178, April 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.27.167;   http://rmp.aps.org/abstract/
RMP/v27/i2/p167_1.

Smoluchowski:1953:I

[Smo53a] R. Smoluchowski. Introduction. Reviews of Modern Physics, 25(1):1, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.25.1;   http://rmp.aps.
org/abstract/RMP/v25/i1/p1_1.

Smoluchowski:1953:MSB

[Smo53b] R. Smoluchowski. Magnetic susceptibility of bromine-graphite. Reviews of Modern Physics, 25(1):178–180, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.178;   http://rmp.aps.org/abstract/RMP/v25/i1/p178_
1.

Smoluchowski:1953:SD

[Smo53c] R. Smoluchowski. Summary of the discussion. Reviews of Modern Physics, 25(1):227–228, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.227;
 http://rmp.aps.org/abstract/RMP/v25/i1/p227_1.

Snow:1953:D

[Sno53a] A. I. Snow. Discussion. Reviews of Modern Physics, 25(1):284, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.284;   http://rmp.aps.org/abstract/
RMP/v25/i1/p284_1.

Snow:1953:MMO

[Sno53b] A. I. Snow. Magnetic moment orientation and thermal expansion of antiferromagnetic CrSb. Reviews of Modern Physics, 25(1):127, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.127;   http://
rmp.aps.org/abstract/RMP/v25/i1/p127_1.

Saenz:1955:NSM

[SO55] A. W. Sáenz and R. C. O’Rourke. Number of states and the magnetic properties of an electron gas. Reviews of Modern Physics, 27(4):381–398, October 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.27.381;
 http://rmp.aps.org/abstract/RMP/v27/i4/p381_1.

Stevens:1953:NEI

[Ste53] K. W. H. Stevens. A note on exchange interactions. Reviews of Modern Physics, 25(1):166, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.166;
 http://rmp.aps.org/abstract/RMP/v25/i1/p166_1.

Stetten:1959:HR

[Ste59] Dewitt Stetten, Jr. Hormone regulation. Reviews of Modern Physics, 31(2):563–568, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.31.563;   http:/
/rmp.aps.org/abstract/RMP/v31/i2/p563_1.

Stoner:1953:AMC

[Sto53] Edmund C. Stoner. The analysis of magnetization curves. Reviews of Modern Physics, 25(1):2–16, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.2;
http://rmp.aps.org/abstract/RMP/v25/i1/p2_1.

Stockmayer:1959:FBM

[Sto59] Walter H. Stockmayer. Forces between macromolecules. Reviews of Modern Physics, 31(1):103–106, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.103;
 http://rmp.aps.org/abstract/RMP/v31/i1/p103_1.

Suess:1956:AE

[SU56] Hans E. Suess and Harold C. Urey. Abundances of the elements. Reviews of Modern Physics, 28(1):53–74, January 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
28.53;   http://rmp.aps.org/abstract/RMP/v28/i1/p53_1.

Sutherland:1959:AIS

[Sut59] G. B. B. M. Sutherland. Application of infrared spectroscopy to biological problems. Reviews of Modern Physics, 31(1):118–122, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.118;   http://rmp.aps.org/abstract/
RMP/v31/i1/p118_1.

Wu:1950:RIS

[sW50] Chien shiung Wu. Recent investigation of the shapes of β-ray spectra. Reviews of Modern Physics, 22(4):386–398, October 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.22.386;   http://rmp.aps.org/abstract/
RMP/v22/i4/p386_1.

Shull:1953:NDS

[SW53] C. G. Shull and M. K. Wilkinson. Neutron diffraction studies of various transition elements. Reviews of Modern Physics, 25 (1):100–107, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.25.100;   http://
rmp.aps.org/abstract/RMP/v25/i1/p100_1.

Tauc:1957:GES

[Tau57] Jan Tauc. Generation of an emf in semiconductors with nonequilibrium current carrier concentrations. Reviews of Modern Physics, 29(3):308–324, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.308;   http://
rmp.aps.org/abstract/RMP/v29/i3/p308_1.

Temmer:1958:ASN

[Tem58] G. M. Temmer. Angular shapes of nuclei. Reviews of Modern Physics, 30(2):498–506, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.498;
 http://rmp.aps.org/abstract/RMP/v30/i2/p498_1.

Tobias:1959:ERP

[Tob59] Cornelius A. Tobias. Effects of radiations on populations of cells and multicellular organisms. Reviews of Modern Physics, 31(2):289–296, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.289;   http://
rmp.aps.org/abstract/RMP/v31/i2/p289_1.

Tolhoek:1956:EPT

[Tol56] H. A. Tolhoek. Electron polarization, theory and experiment. Reviews of Modern Physics, 28(3):277–298, July 1956. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.28.277;
 http://rmp.aps.org/abstract/RMP/v28/i3/p277_1.

Uttley:1959:IPR

[Utt59] Albert M. Uttley. Imitation of pattern recognition and trial-and-error learning in a conditional probability computer. Reviews of Modern Physics, 31(2):546–548, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.546;
 http://rmp.aps.org/abstract/RMP/v31/i2/p546_1.

VanVleck:1951:CAM

[Van51] J. H. Van Vleck. The coupling of angular momentum vectors in molecules. Reviews of Modern Physics, 23(3):213–227, July 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.23.213;   http://rmp.aps.org/abstract/
RMP/v23/i3/p213_1.

VanVleck:1953:MEC

[Van53] J. H. Van Vleck. Models of exchange coupling in ferromagnetic media. Reviews of Modern Physics, 25(1):220–227, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.220;   http://rmp.aps.org/abstract/
RMP/v25/i1/p220_1.

VanHove:1957:SMS

[Van57] Léon Van Hove. Statistical mechanics: A survey of recent lines of investigation. Reviews of Modern Physics, 29(2):200–204, April 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.200;   http://rmp.aps.org/abstract/
RMP/v29/i2/p200_1.

Varshni:1957:CSP

[Var57] Yatendra Pal Varshni. Comparative study of potential energy functions for diatomic molecules. Reviews of Modern Physics, 29(4):664–682, October 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.664;   http://
rmp.aps.org/abstract/RMP/v29/i4/p664_1. See [Var59].

Varshni:1959:ECS

[Var59] Yatendra Pal Varshni. Errata: Comparative Study of Potential Energy Functions for Diatomic Molecules. Reviews of Modern Physics, 31(3):839, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.839;   http://
rmp.aps.org/abstract/RMP/v31/i3/p839_1. See [Var57].

vandeHulst:1958:DVD

[vdH58] H. C. van de Hulst. Density and velocity distribution of the interstellar gas. Reviews of Modern Physics, 30(3):913–923, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.30.913;   http://rmp.aps.org/abstract/
RMP/v30/i3/p913_1.

Verma:1959:UAD

[Ver59] G. S. Verma. Ultrasonic absorption due to chemical relaxation in electrolytes. Reviews of Modern Physics, 31(4):

1052–1071, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.31.1052;   http://
rmp.aps.org/abstract/RMP/v31/i4/p1052_1.

vonHippel:1950:FDS

[vH50] A. von Hippel. Ferroelectricity, domain structure, and phase transitions of barium titanate. Reviews of Modern Physics, 22(3):221–237, July 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.22.221;   http://
rmp.aps.org/abstract/RMP/v22/i3/p221_1.

VanKranendonk:1958:SW

[VV58] J. Van Kranendonk and J. H. Van Vleck. Spin waves. Reviews of Modern Physics, 30(1):1–23, January 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.1;
http://rmp.aps.org/abstract/RMP/v30/i1/p1_1. See errata

[VV59].

VanKranendonk:1959:ESW

[VV59] J. Van Kranendonk and J. H. Van Vleck. Errata: Spin Waves. Reviews of Modern Physics, 31(3):839, July 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
31.839;   http://rmp.aps.org/abstract/RMP/v31/i3/p839_

3. See [VV58].

VanPatter:1954:NDE

[VW54] D. M. Van Patter and Ward Whaling. Nuclear disintegration energies. Reviews of Modern Physics, 26(4):402–443, October 1954. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.26.402;   http://rmp.aps.org/abstract/
RMP/v26/i4/p402_1.

VanPatter:1957:NDE

[VW57] D. M. Van Patter and Ward Whaling. Nuclear disintegration energies. II. Reviews of Modern Physics, 29(4):757–766, October 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.29.757;   http://rmp.aps.org/abstract/
RMP/v29/i4/p757_1.

Watanabe:1955:SPLa

[Wat55a] Satosi Watanabe. Symmetry of physical laws Part I. Symmetry in space-time and balance theorems. Reviews of Modern Physics, 27(1):26–39, January 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.27.26;   http://
rmp.aps.org/abstract/RMP/v27/i1/p26_1.

Watanabe:1955:SPLb

[Wat55b] Satosi Watanabe. Symmetry of physical laws Part II. Q-number theory of space-time inversions and charge conjugation. Reviews of Modern Physics, 27(1):40–76, January 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
27.40;   http://rmp.aps.org/abstract/RMP/v27/i1/p40_1.

Watanabe:1955:SPLc

[Wat55c] Satosi Watanabe. Symmetry of physical laws. Part III. Prediction and retrodiction. Reviews of Modern Physics, 27(2):179–186, April 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.27.179;   http://rmp.aps.org/abstract/
RMP/v27/i2/p179_1.

Watson:1958:IDI

[Wat58] Kenneth M. Watson. Implications of the direct-interaction model for nuclear structure. Reviews of Modern Physics, 30 (2):565–568, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.565;   http://
rmp.aps.org/abstract/RMP/v30/i2/p565_1.

Waugh:1959:BCS

[Wau59a] David F. Waugh. Blood coagulation — a study in homeostasis. Reviews of Modern Physics, 31(2):557–562, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.557;
 http://rmp.aps.org/abstract/RMP/v31/i2/p557_1.

Waugh:1959:PTI

[Wau59b] David F. Waugh. Proteins and their interactions. Reviews of Modern Physics, 31(1):84–93, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
31.84;   http://rmp.aps.org/abstract/RMP/v31/i1/p84_1.

Williams:1953:MSL

[WB53] H. J. Williams and R. M. Bozorth. Magnetic study of low temperature transformation in magnetite. Reviews of Modern Physics, 25(1):79–80, January 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.25.79;   http://
rmp.aps.org/abstract/RMP/v25/i1/p79_1.

Weiss:1958:XRD

[WD58] R. J. Weiss and J. J. Demarco. X-ray determination of the number of 3d electrons in Cu, Ni, Co, Fe, and Cr. Reviews of Modern Physics, 30(1):59–62, January 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
30.59;   http://rmp.aps.org/abstract/RMP/v30/i1/p59_1.

Weber:1959:M

[Web59] J. Weber. Masers. Reviews of Modern Physics, 31(3):681–710, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.681;   http://rmp.aps.org/abstract/

RMP/v31/i3/p681_1. See erratum [Web60].

Weber:1960:EM

[Web60] J. Weber. Erratum: Masers. Reviews of Modern Physics, 32(4):1033, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.1033;   http://
rmp.aps.org/abstract/RMP/v32/i4/p1033_2. See [Web59].

Weil:1953:TFF

[Wei53] L. Weil. The texture of fine ferromagnetic powders. Reviews of Modern Physics, 25(1):324–326, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.324;
 http://rmp.aps.org/abstract/RMP/v25/i1/p324_1.

Weisskopf:1957:NP

[Wei57] Victor F. Weisskopf. Nuclear physics. Reviews of Modern Physics, 29(2):174–181, April 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.174;   http://
rmp.aps.org/abstract/RMP/v29/i2/p174_1.

Weiss:1959:CD

[Wei59a] Paul Weiss. Cellular dynamics. Reviews of Modern Physics, 31(1):11–20, January 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.31.11;   http://
rmp.aps.org/abstract/RMP/v31/i1/p11_1.

Weiss:1959:IBC

[Wei59b] Paul Weiss. Interactions between cells. Reviews of Modern Physics, 31(2):449–454, April 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.31.449;   http://
rmp.aps.org/abstract/RMP/v31/i2/p449_1.

Wexler:1958:DAB

[Wex58] Sol Wexler. Deposition of atomic beams. Reviews of Modern Physics, 30(2):402–409, April 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.402;   http:/
/rmp.aps.org/abstract/RMP/v30/i2/p402_1.

Wijn:1953:NHF

[WGvdB53] H. P. J. Wijn, M. Gevers, and C. M. van der Burgt. Note on the high frequency dispersion in nickel zinc ferrites. Reviews of Modern Physics, 25(1):91–92, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.91;   http://rmp.aps.org/abstract/RMP/v25/i1/p91_1.

Wheeler:1957:AER

[Whe57] John A. Wheeler. Assessment of Everett’s “relative state” formulation of quantum theory. Reviews of Modern Physics, 29(3):463–465, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.463;   http://
rmp.aps.org/abstract/RMP/v29/i3/p463_1. See [Eve57].

White:1955:MHS

[Whi55] Robert L. White. Magnetic hyperfine structure due to rotation in ^{1}Σ molecules. Reviews of Modern Physics, 27(3):276–288, July 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.27.276;   http://rmp.aps.org/abstract/
RMP/v27/i3/p276_1.

Whipple:1958:OA

[Whi58] F. L. Whipple. Opening address. Reviews of Modern Physics, 30(3):912, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.30.912;   http://
rmp.aps.org/abstract/RMP/v30/i3/p912_1.

Wick:1955:ISR

[Wic55] G. C. Wick. Introduction to some recent work in meson theory. Reviews of Modern Physics, 27(4):339–362, October 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
27.339;   http://rmp.aps.org/abstract/RMP/v27/i4/p339_
1.

Wigner:1957:RIQ

[Wig57] Eugene P. Wigner. Relativistic invariance and quantum phenomena. Reviews of Modern Physics, 29(3):255–268, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.255;   http://rmp.aps.org/abstract/
RMP/v29/i3/p255_1.

Wilets:1958:TNS

[Wil58a] Lawrence Wilets. Theories of the nuclear surface. Reviews of Modern Physics, 30(2):542–549, April 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.542;
 http://rmp.aps.org/abstract/RMP/v30/i2/p542_1.

Wilson:1958:IKP

[Wil58b] O. C. Wilson. Internal kinematics of the planetary nebulae. Reviews of Modern Physics, 30(3):1025–1029, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.30.1025;   http://rmp.aps.org/abstract/RMP/
v30/i3/p1025_1.

Wilcox:1959:RHT

[Wil59a] John M. Wilcox. Review of high-temperature rotating-plasma experiments. Reviews of Modern Physics, 31(4):1045–1051, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.31.1045;   http://rmp.aps.org/abstract/

RMP/v31/i4/p1045_1. See comment [Sle60].

Williams:1959:RNA

[Wil59b] Robley C. Williams. Replication of nucleic acids. Reviews of Modern Physics, 31(1):233–241, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.233;
 http://rmp.aps.org/abstract/RMP/v31/i1/p233_1.

Wilcox:1960:RCR

[Wil60] John M. Wilcox. Reply to comment on “Review of High-Temperature Rotating-Plasma Experiments”. Reviews of Modern Physics, 32(4):1033, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.1033;   http:
//rmp.aps.org/abstract/RMP/v32/i4/p1033_1. See [Wil59a,

Sle60].

Wohlfarth:1953:TES

[Woh53] E. P. Wohlfarth. The theoretical and experimental status of the collective electron theory of ferromagnetism. Reviews of Modern Physics, 25(1):211–219, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.211;
 http://rmp.aps.org/abstract/RMP/v25/i1/p211_1.

Wood:1959:SAC

[Woo59] T. H. Wood. Some aspects of cellular radiobiology. Reviews of Modern Physics, 31(2):282–288, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.282;
 http://rmp.aps.org/abstract/RMP/v31/i2/p282_1.

Waser:1953:FID

[WS53] Jürg Waser and Verner Schomaker. The Fourier inversion of diffraction data. Reviews of Modern Physics, 25(3):671–690, July 1953. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.25.671;   http://rmp.aps.org/abstract/
RMP/v25/i3/p671_1.

Wu:1959:PEB

[Wu59] C. S. Wu. Parity experiments in beta decays. Reviews of Modern Physics, 31(3):783–790, July 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.783;
 http://rmp.aps.org/abstract/RMP/v31/i3/p783_1.

Wijn:1953:RTA

[WvdH53] H. P. J. Wijn and H. van der Heide. A richter type aftereffect in ferrites containing ferrous and ferric ions. Reviews of Modern Physics, 25(1):98–99, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
25.98;   http://rmp.aps.org/abstract/RMP/v25/i1/p98_1.

Weber:1957:RCG

[WW57] Joseph Weber and John A. Wheeler. Reality of the cylindrical gravitational waves of Einstein and rosen. Reviews of Modern Physics, 29(3):509–515, July 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.509;
 http://rmp.aps.org/abstract/RMP/v29/i3/p509_1.

Witte:1958:EDN

[WW58] H. Witte and E. Wölfel. Electron distributions in NaCl, LiF, CaF_{2}, and Al. Reviews of Modern Physics, 30(1): 51–55, January 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.51;   http://
rmp.aps.org/abstract/RMP/v30/i1/p51_1.

Yang:1957:PKA

[Yan57] C. N. Yang. Present knowledge about the new particles. Reviews of Modern Physics, 29(2):231–235, April 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.231;
 http://rmp.aps.org/abstract/RMP/v29/i2/p231_1.

Chen:1957:BSS

[yCT57] Shang yi Ch’en and Makoto Takeo. Broadening and shift of spectral lines due to the presence of foreign gases. Reviews of Modern Physics, 29(1):20–73, January 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
29.20;   http://rmp.aps.org/abstract/RMP/v29/i1/p20_1.

Yvon:1953:ECB

[YHA53] J. Yvon, J. Horowitz, and A. Abragam. Exchange coupling between three iron ions in two organic molecules. Reviews of Modern Physics, 25(1):165, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.165;
 http://rmp.aps.org/abstract/RMP/v25/i1/p165_1.

Yennie:1957:ESN

[YLR57] D. R. Yennie, M. M. Lévy, and D. G. Ravenhall. Electromagnetic structure of nucleons. Reviews of Modern Physics, 29 (1):144–157, January 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.29.144;   http://
rmp.aps.org/abstract/RMP/v29/i1/p144_1.

Yukawa:1957:IRM

[Yuk57] Hideki Yukawa. Introductory remarks on meson theory. Reviews of Modern Physics, 29(2):213–215, April 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.213;
 http://rmp.aps.org/abstract/RMP/v29/i2/p213_1.

Zanstra:1958:CGR

[Zan58] H. Zanstra. Comparison of gas and radiation pressure in the problems of the dynamics of planetary nebulae. Reviews of Modern Physics, 30(3):1030–1033, July 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.30.1030;   http:
//rmp.aps.org/abstract/RMP/v30/i3/p1030_1.

Zener:1953:EI

[ZH53] C. Zener and R. R. Heikes. Exchange interactions. Reviews of Modern Physics, 25(1):191–198, January 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.191;
 http://rmp.aps.org/abstract/RMP/v25/i1/p191_2.

Zimm:1959:CMS

[Zim59] Bruno H. Zimm. Concentrated macromolecular solutions. Reviews of Modern Physics, 31(1):123–129, January 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.123;
 http://rmp.aps.org/abstract/RMP/v31/i1/p123_1.

Zinn:1955:FAE

[Zin55] Walter H. Zinn. Fermi and atomic energy. Reviews of Modern Physics, 27(3):263–268, July 1955. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.27.263;
 http://rmp.aps.org/abstract/RMP/v27/i3/p263_1.

Zirkle:1959:GFR

[Zir59] Raymond E. Zirkle. General features of radiobiological actions. Reviews of Modern Physics, 31(2):269–272, April 1959. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.31.269;
 http://rmp.aps.org/abstract/RMP/v31/i2/p269_1.